### v1

# Towards Principled Assessment of Tabular Data Synthesis Algorithms

Abstract
Data synthesis has been advocated as an important approach for utilizing data while protecting data privacy. A large number of tabular data synthesis algorithms (which we call synthesizers) have been proposed. Some synthesizers satisfy Differential Privacy, while others aim to provide privacy in a heuristic fashion. A comprehensive understanding of the strengths and weaknesses of these synthesizers remains elusive due to lacking principled evaluation metrics and missing head-to-head comparisons of newly developed synthesizers that take advantage of diffusion models and large language models with state-of-the-art marginal-based synthesizers.
In this paper, we present a principled and systematic evaluation framework for assessing tabular data synthesis algorithms. Specifically, we examine and critique existing evaluation metrics, and introduce a set of new metrics in terms of fidelity, privacy, and utility to address their limitations. Based on the proposed metrics, we also devise a unified objective for tuning, which can consistently improve the quality of synthetic data for all methods. We conducted extensive evaluations of 8 different types of synthesizers on 12 datasets and identified some interesting findings, which offer new directions for privacy-preserving data synthesis.
1 Introduction
Data-driven decision-making has emerged as the prevailing approach to advance science, industrial applications, and governance, creating the necessity to share and publish tabular data. At the same time, growing concerns about the privacy breaches caused by data disclosure call for data publishing approaches that preserve privacy. One increasingly advocated and adopted approach to reduce privacy risks while sharing data is to release synthetic data. Ideally, synthetic data can effectively fit any data processing workflow designed for the original data without privacy concerns. Data synthesis initiatives have been promoted not only by the research community [58, 29] but also among non-profit organizations [45, 61] and government agencies [5, 43, 44].
In this paper, we study data synthesis algorithms for tabular data, which we call synthesizers. In recent years, a plethora of synthesizers have been proposed, which can be roughly categorized into two groups: statistical and deep generative. Statistical synthesizers [69, 41, 39, 70] use low-order marginals of the real data to create synthetic datasets that match real distributions. They were the best-performing algorithms in NIST competitions [43, 44]. Deep generative synthesizers [48, 30, 65, 32, 7], on the other hand, learn the data distribution from real data records and generate synthetic instances by sampling from the distribution. With the recent development in deep generative models (e.g., diffusion models [26] and large language models [64, 54]), new synthesizers were proposed [32, 7, 31, 68] to extend these successes to the realm of tabular data synthesis.
While recent state-of-the-art approaches achieve compelling results in synthesizing authentic data, a comprehensive understanding of the strengths and weaknesses of different synthesizers remains elusive. We are unaware of any head-to-head comparison of the best statistical synthesizers [41, 39, 70] against the best deep generative synthesizers [32, 7, 31].
In addition, there is a lack of principled and widely accepted evaluation metrics. It is known that evaluating synthesizers is inherently difficult [59, 12, 2, 29], and qualitative evaluation of tabular data through visual inspection is also infeasible. Table 1 shows that existing papers use different evaluation metrics, resulting in substantially disparate conclusions.
Table 1: Main evaluation protocols for synthesizers in the literature. “-” means that the synthesizer does not consider this property. “ ✓ ” denotes that the model is viewed as the state-of-the-art method by many related studies [40, 20, 46, 58]. “ ✓   ” denotes that the model is frequently used in recent papers[32, 34, 31] for comparison as a state-of-the-art method.
Algorithm Privacy Type Architecture Fidelity Utility State-of-the-art
PGM [39, 41] Differentially Private Probabilistic Graphical Model Total Variation Distance Range Query
PrivSyn [70] Differentially Private Non-parametric Model Total Variation Distance Machine Learning Efficacy + Range Query
TVAE [65] Heuristically Private Variational Autoencoder Likelihood Fitness Machine Learning Efficacy
CTGAN [65] Heuristically Private Generative Adversarial Network Likelihood Fitness Machine Learning Efficacy
PATE-GAN [30] Differentially Private Generative Adversarial Network - Machine Learning Efficacy
TabDDPM [32] Differentially Private Diffusion Model Correlations Machine Learning Efficacy
TableDiffusion [60] Differentially Private Diffusion Model Kolmogorov-Smirnov Test Machine Learning Efficacy
GReaT [7] Heuristically Private Large Language Model - Machine Learning Efficacy
Since synthetic data is proposed to protect the privacy of real data, privacy should be an indispensable part of the evaluation. While many differentially private (DP) synthesizers [69, 39, 70, 40] have been proposed with provable privacy guarantees, they often suffer from significant utility loss [58, 20, 27]. Therefore, a parallel line of research focuses on what we call heuristically private (HP) synthesizers [48, 71, 32, 68, 7, 31], which do not integrate DP for model learning. These methods typically rely on similarity-based metrics to empirically evaluate privacy risks [^1] These metrics are also widely used in industry, see [19] for detailed analysis.. The underlying rationale is that the closer the synthetic data is to real data, the higher the risk of information leakage. However, the meaning of these metrics has not been sufficiently examined. Ironically, recent attacks [19] utilize the DCR (one of the most popular similarity-based privacy metrics, detailed in Section 4.1) to successfully recover a large proportion of outliers in real data, casting doubt on the effectiveness and reliability of current privacy measurements.
Another problem of existing evaluations is the prevalent reliance on default hyperparameter settings for training synthesizers [57, 7, 68, 31]. This approach can result in suboptimal performance of synthesizers and skewed comparisons, leading to misleading conclusions (we summarize them in Appendix E). The root of this problem lies in the fact that many synthesizers do not provide details about hyperparameter selections, posing a challenge in conducting fair comparisons among diverse synthesizers on an equal footing.
The above concerns motivate us to design a principled evaluation framework for data synthesis, aimed at elucidating the current advancements in this field. We believe that principled and systematic assessments are crucial: not only can they help practitioners choose from a large set of synthesizers, but they can also aid in understanding the strengths and weaknesses of each type of method, making progress toward developing better synthesis algorithms.
1.1 Our Contributions
In this paper, we examine, characterize, and critique the commonly used evaluation metrics, and propose a principled framework as well as a set of new metrics for data synthesis evaluation. Our assessments unfold along three main axes:
  • Fidelity. We identify faithfulness and universality as two key properties for fidelity metrics in Section 2.1 and find that none of the existing fidelity metrics can satisfy both. To address this gap, we introduce a new fidelity metric based on the Wasserstein distance, which provides a unified way to evaluate the numerical, discrete, and mixture attributes under the same criteria.
  • Privacy. We address the inadequacy of current similarity-based privacy metrics and membership inference (MI) attacks and propose the novel membership disclosure score to directly gauge the disclosure risk of synthesizers without relying on membership inference (MI) attacks. Our measurement is aligned with the principles of DP and only requires black-box access to synthesizers.
  • Utility. We advocate using two tasks for assessing the utility of synthetic data: machine learning prediction and range (point) query. To eliminate the inconsistent performance caused by the choice of different machine learning models, we introduce a new and robust metric to quantify the distribution shift of synthetic data.
SynMeter. We implement a systematic evaluation framework called SynMeter to support the assessment of data synthesis algorithms with the proposed metrics. SynMeter differs from previous approaches in that, instead of relying on default hyperparameters of different synthesizers, it incorporates a model tuning phase. SynMeter combines the proposed metrics to create a novel and unified objective for model tuning, which eases hyperparameter selection and consistently improves the quality of synthetic data for all methods. With a modular design, SynMeter can easily integrate any additional synthesizers and datasets by implementing new functional codes to the relevant modules. Our code is publicly available [^2] https://github.com/zealscott/SynMeter, facilitating researchers to leverage existing pipelines to tune, assess, or benchmark (new) data synthesis algorithms.
Main Findings. Extensive experiments have been conducted on a wide range of state-of-the-art heuristically private (HP) and differentially private (DP) synthesis algorithms (list in Table 1) over 12 real-world datasets. Our evaluation finds that no single synthesis algorithm emerges as the definitive leader. Synthesizers struggle to balance competitive performance with robust privacy measures, exhibiting a pronounced performance disparity between HP and DP synthesis algorithms. Specifically, we have the following interesting findings:
  • Diffusion models are surprisingly good at synthesizing tabular data. Although they were originally introduced for image generation, our evaluation indicates their expressive capability of synthesizing highly authentic tabular data. For instance, the diffusion-based synthesizer (i.e., TabDDPM) surpasses its counterparts in terms of both fidelity and utility. However, it suffers from significant membership privacy risks, and directly applying differential privacy to it would negate its advantages, highlighting a critical challenge for further exploration.
  • Statistical methods are still competitive synthesizers. State-of-the-art statistical methods consistently outperform all deep generative synthesizers when DP is required. Impressively, even in heuristically private settings, these statistical approaches still maintain competitive performance over many sophisticated synthesizers like TVAE and GReaT.
  • Disparate impacts of privacy budget on DP synthesizers: We observe that statistical methods like PGM can achieve good performance even with a small privacy budget (e.g., $\epsilon=0.5)$. However, deep generative synthesizers generally require a much larger privacy budget (e.g., $\epsilon=8)$ to yield comparable results.
  • Large language models (LLM) are semantic-aware synthesizers. LLM-based method (i.e., GReaT) excels at generating realistic tabular data when the input dataset consists of rich semantic attributes, establishing a new paradigm in data synthesis.
  • Despite CTGAN being widely recognized as the strong baseline in data synthesis, a closer examination using proposed metrics revealed that it struggles to learn marginal distributions during training, which leads to unsatisfactory results on complex tabular data.
Related Studies. Several studies [58, 27, 3, 9] have been conducted to benchmark tabular synthesis algorithms. However, they all focus on DP synthesizers, neglecting extensively studied HP synthesizers. In particular, they did not include the newly proposed deep generative synthesizers using diffusion models or large language models [32, 7, 31], which we found to be very promising in our studies. Furthermore, they directly use existing metrics for comparisons, while we identify the limitations of these metrics and propose a new and more principled assessment for tabular data synthesis.
Roadmap. We first propose a principled framework for tabular data synthesis evaluation in Section 2. Then from Section 3 to Section 5, we discuss existing evaluation methods and propose our metrics from fidelity, privacy and utility aspects. In Section 6 we introduce a unified tuning objective as well as our framework implementation. The experimental results are shown in Section 7. Finally, we summarize our findings and conclude the paper in Section 8.
2 A Framework for Data Synthesis Evaluation
In this section, we introduce a framework for evaluating the effectiveness of synthesizers.
2.1 Evaluation Criteria
A synthesizer $\mathsf{A}$ takes as input a dataset $\mathcal{D}$, which is assumed to be sampled i.i.d. from some (often unknown) underlying data distribution $\mathbb{D}$, and outputs a generator $\mathsf{G}$. $\mathsf{G}$ on input an integer $n$ outputs a dataset $\mathcal{S}$ with $n$ data instances. We consider three classes of desirable properties for data synthesis: Fidelity, Privacy, and Utility.
Fidelity. We want the distribution for the synthetic data instances to be close to $\mathbb{D}$. Since $\mathbb{D}$ is often unknown, fidelity is measured by the distance between the input dataset $\mathcal{D}$ and the synthesized dataset $\mathcal{S}$. If one partitions the input dataset $\mathcal{D}$ into a training set $\mathcal{D}_{\text{train}}$ and a testing set $\mathcal{D}_{\text{test}}$, one can measure fidelity as closeness to either $\mathcal{D}_{\text{train}}$ or $\mathcal{D}_{\text{test}}$. Given the heterogeneity and high-dimensional nature of tabular data, assessing the distribution similarity can be challenging. We use the following desiderata for fidelity metrics:
  • Faithfulness. An effective fidelity metric should faithfully reveal the difference between the distributions of synthetic data and real data.
  • Universality. To tackle the heterogeneity of tabular data, the metric should accommodate both numerical and categorical attributions, and avoid reliance on a plethora of ad-hoc measures for different data types.
Privacy. Using synthetic data is usually motivated by the desire to protect the input dataset. Without this consideration, one should share the input dataset to achieve the best fidelity and utility. For privacy, differential privacy [16] is generally considered to be the golden standard. Some synthesizers are designed to satisfy DP, either by directly perturbing data statistics or adding noises during training. However, satisfying DP under reasonable parameters may result in poor performance; heuristically private (HP) synthesizers do not satisfy DP, and aim to protect privacy empirically. We thus need empirical metrics for privacy evaluations, and the following requirements should be considered:
  • A privacy metric should reflect the foundational concept of privacy, i.e., the ability to protect every training instance against membership disclosure, as it is well understood that DP is equivalent to membership privacy when membership of different instances are independent [35]. Heuristic approaches [71] such as the closeness to real data deviate from this goal and violate this requirement.
  • A reliable privacy metric should be able to differentiate the membership disclosure risks among various synthesizers. While membership inference (MI) attacks against classifiers have been extensively studied, MI attacks against synthesizers either rely on unrealistic settings [57] or are too weak for evaluations [28]. Using MI attacks to evaluate the privacy of synthesizers is thus problematic.
We note that privacy metrics are also useful for evaluating DP synthesizers, as the privacy analysis may not be tight and the degrees of tightness vary based on the algorithms. Two algorithms with the same DP parameters $\epsilon,\delta$ may be offering different levels of privacy protection (detailed in Section 7.6).
Utility. Sometimes, when using synthesizers, one know the specific tasks the synthesized datasets will be used for, and desires that the synthesized dataset $\mathcal{S}$ is useful for tasks at hand, which may be achievable without high fidelity. We consider the following requirements for utility metrics that measure effectiveness of synthesized datasets for specific tasks.
  • The tasks chosen for evaluating synthetic data should be both representative and well-defined.
  • The performance on these tasks should be consistent and reliable. Since some tasks use auxiliary evaluators (e.g., machine learning models), it is essential to ensure that the choice of these evaluators does not affect the evaluations.
2.2 Evaluation Pipelines
We propose SynMeter, a systematic evaluation framework for synthesizers. SynMeter consists of four phases: data preparation, model tuning, model training, and model evaluation.
The data preparation phase explores and preprocesses data for learning algorithms [^3] Here we assume that there are no missing values in original data. The missing values problem has been extensively studied [51], which is independent of data synthesis.. In this phase, statistical methods [69, 70, 39] select low-dimensional marginals to serve as a compact representation for capturing data dependencies. Deep generative models [65, 26, 68, 71] involve standard data processing techniques like data encoding and normalization.
The goal of model tuning phase is to select optimal hyperparameters for generative models. A subset of training dataset $\mathcal{D}_{\text{train}}$ is sampled as validation data $\mathcal{D}_{\text{val}}$. The generative model $\mathsf{G}$ is trained on $\mathcal{D}_{\text{train}}$ and assessed its performance on $\mathcal{D}_{\text{val}}$ for model tuning. Note that the tuning phase is often overlooked in many data synthesis models. However, we show in Section 7.6 that a proper tuning objective can boost the performance of synthesizers.
The model training phase focuses on model learning with tuned hyperparameters. Various generative models implement different architectures and optimization objectives.
In the model evaluation phase, the trained model generates some synthetic data $\mathcal{S}$, which are used for evaluation.
3 Fidelity Evaluation
In this section, we first review existing fidelity measurements and identify their limitations, and then introduce a new and more principled fidelity metric.
3.1 Existing Metrics and Limitations
Low-order Statistics. Low-dimensional marginals are crucial to the quality of synthetic data and many well-established statistics are used for one (two)-way marginals evaluation:
Distribution Measurements. Total Variation Distance (TVD) [70, 39, 32] and Kolmogorov-Smirnov Test (KST) [68] are used to measure the univariate distribution similarity for categorical and numerical attributes, respectively. TVD computes the absolute probabilities difference of categorical distributions, while KST examines the maximum discrepancy between cumulative distribution functions for numerical attributes. The main problem of these methods is the lack of versatility. Each type of marginal requires a distinct statistical measure, which complicates the ability to perform a principled comparison across various data types.
Correlation Statistics. Studies [12, 32, 71, 58] often utilize correlation differences for pairwise similarity evaluations. Specifically, the correlation score of any two pairs of attributes is calculated on both synthetic and real data, and the overall mean difference of correlation scores is used as fidelity. Popular correlation statistics like Theil’s uncertainty coefficient [^4] https://en.wikipedia.org/wiki/Uncertainty_coefficient, Pearson correlation, and the correlation ratio [^5] https://en.wikipedia.org/wiki/Correlation_ratio are applied for various marginal types (categorical, continuous, and mixed). However, correlation coefficients are scale-invariant, meaning they do not take into account differences in the scale (variance) or location (mean) of distributions, which would result in a misleading perception of similarity between the synthetic and real datasets.
Likelihood Fitness. This metric [65] aims to directly compute the likelihood of synthetic data as fidelity by using simulated data as input datasets. Since the input data is generated from some prior probabilistic models (e.g., Bayesian networks), the likelihood of synthetic data can be derived by fitting them to the priors. However, this approach can only be applied by simulation, which limits the application for real-world datasets.
Table 2: Comparison of existing fidelity metrics. “ ✓ ” indicates the metric meets the corresponding property, and “ ✗ ” denotes otherwise.
Fidelity Metric Type Faithfulness Universality
TVD/KST Statistics
Correlations Statistics
Likelihood Fitness Simulation
$\alpha$-Precision Coverage-based
Probabilistic MSE Model-based
Wasserstein Distance (ours) Statistics
Other Approaches. $\alpha$-Precision [2] defines fidelity as the proportion that the generated sample is covered by real data. The coverage is determined by comparing the distance between synthetic data and the nearest real data to a target real one. Probabilistic mean squared error (pMSE) [56] employs a logistic regression discriminator to distinguish between synthetic and real data, using relative prediction confidence as a fidelity metric. However, these heuristic approaches are unable to accurately capture the distribution difference between synthetic data and real data.
Discussion. We compare existing metrics against the principles of fidelity evaluation, as outlined in Table 2. The results indicate that none of these metrics fulfill all the desired properties. This gap underlines the necessity for developing a new fidelity metric for holistic assessment.
3.2 Proposed Metric: Wasserstein Distance
We opt for Wasserstein distance to measure the distribution discrepancies between synthetic data and real data. Originating from optimal transport theory [50], the Wasserstein distance provides a structure-aware measure of the minimal amount of work required to transform one distribution into another. Formally, Let $\mathbf{P}=(p_{1},p_{2},\ldots p_{n})$ and $\mathbf{Q}=(q_{1},q_{2},\ldots q_{n})$ be the two probability distributions, and $\mathbf{C}$ be a matrix of size $n\times n$ in which $\mathbf{C}_{ij}\geq 0$ is the cost of moving an element $i$ of $\mathbf{P}$ to the element $j$ of $\mathbf{Q}$ ($\mathbf{C}_{ii}=0$ for all element $i$). The optimal transport plan $\mathbf{A}$ is defined as the minimal amount of work between two distributions:
$\begin{split}&\min_{\mathbf{A}}\quad\langle\mathbf{C},\mathbf{A}\rangle\\ \text{ s.t. }&\mathbf{A}\mathds{1}=\mathbf{P},\quad\mathbf{A}^{\top}\mathds{1}=\mathbf{Q}\end{split}$ (1)
[…]
$\mathcal{W}(\mathbf{P},\mathbf{Q})=\langle\mathbf{C},\mathbf{A}^{*}\rangle$ (2)
Now we can use Wasserstein distance to define the fidelity:
Definition 1
(Wasserstein-based Fidelity Metric) Let $\mathbf{P}$ and $\mathbf{Q}$ be the marginal distributions derived from synthetic data $\mathcal{S}$ and real data $\mathcal{D}$, the fidelity of data synthesis algorithm $\mathsf{A}$ is:
$\mathrm{Fidelity}(\mathsf{A})\coloneqq\underset{\begin{subarray}{c}\mathbf{P}\sim\mathcal{S}\\ \mathbf{Q}\sim\mathcal{D}\end{subarray}}{\mathbb{E}}\big{[}\mathcal{W}(\mathbf{P},\mathbf{Q})\big{]}$ (3)
where $\mathcal{W}(\mathbf{P},\mathbf{Q})$ is the Wasserstein distance between synthetic and real marginal distributions.
In practice, the dataset $\mathcal{D}$ can be designated as either $\mathcal{D}_{train}$ or $\mathcal{D}_{test}$, depending on the specific requirements of the analysis. We compute the Wasserstein distance of all the one-way and two-way marginals and use the mean as the final fidelity score.
Determining Cost Matrix. The Wasserstein distance requires the predefined cost matrix $\mathbf{C}$, which encapsulates the “difficulty” or “cost” of transitioning from one distribution element to another. For $k$-way marginal distributions $\mathbf{P}$ and $\mathbf{Q}$, the cost matrix is formulated by summing the pairwise distances between corresponding elements:
$\mathbf{C}_{ij}=\sum\nolimits_{r}^{k}d(v_{i}^{r},v_{j}^{r}).$ (4)
Here, $v_{i},v_{j}\in\mathbb{R}^{k}$ are the element located in $i$ and $j$ in $k$-way probability distributions, respectively. The distance $d(\cdot,\cdot)$ is tailored to the nature of the elements, differing for numerical and categorical values:
$d(v_{i}^{r},v_{j}^{r})=\begin{cases}||v_{i}^{r}-v_{j}^{r}||_{1}&\text{if numerical}\\ \infty,&\forall v_{i}^{r}\neq v_{j}^{r},\text{if categorical}.\end{cases}$ (5)
This means we use $l_{1}$ distance for numerical values and consider the strict match for categorical values (i.e., cost for any different categorical values is infinity).
Solving Wasserstein Distance. The computation of Wasserstein distance involves solving the optimization problem in Equation 1, which is an instance of linear programming problems (LP) [50]. There are many open-source libraries like CVXPY [14] and OPT [18] that can be used to solve LP reasonably fast. Besides, it can also be computed by solving an instance of minimum cost flow problem using the network simplex algorithm, which is much faster than using the LP solver when the cost matrix $\mathbf{C}$ is sparse [36].
However, when the cost matrix is large and dense, computing this metric can be prohibitively time-consuming. Several options are provided to address this problem. (i) Sinkhorn distance [13] provides a fast approximation to the Wasserstein distance by penalizing the objective with an entropy term. (ii) Sliced-Wasserstein distance [6], which uses Radon transform to linearly project data into one dimension, can be efficiently computed. (iii) Reducing the size of the cost matrix by randomly sampling a small set of points from the probability densities. These computational strategies ensure the Wasserstein distance remains a feasible solution for fidelity evaluation, even for complex and high-dimensional marginals.
Advantages of Wasserstein-based Fidelity Metric. Wasserstein distance offers several distinct advantages for fidelity evaluations: (i) It serves as a natural and structure-aware measurement for analyzing distribution discrepancies, which generalizes existing metrics such as TVD and contingency similarity [49, 68]. (ii) It can accommodate both numerical and categorical attributes under the same criterion, which facilitates the evaluation of heterogeneous types of marginals.
4 Privacy Evaluation
In this section, we examine the popular privacy evaluation methodologies employed for synthesizers and propose a novel and effective privacy metric: membership disclosure score.
4.1 Existing Metrics and Limitations
Similarity-based Privacy Metrics. Similarity-based metrics have become the conventional privacy evaluation metrics for HP synthesizers [38, 19, 29]. These heuristic methods empirically assess the privacy risks by measuring the closeness between synthetic data and real data. Among all heuristics, the Distance to Closest Records (DCR) is the most widely used both in academia [71, 66, 32, 68] and industry [4, 23, 42]. DCR looks at the distribution of the distances from each synthetic data point to its nearest real one, and uses the 5th percentile (or the mean) of this distribution as the privacy score. A small score indicates that the synthetic dataset is too close (similar) to real data, signaling a high risk of information leakage.
Although the prevalence of these metrics as privacy evaluation, the meaningfulness of these heuristics has never been rigorously studied. The commonly accepted concept of privacy breach is the ability of an adversary to either re-identify or assert the membership of an individual in real data [35, 16, 61, 45]. Hence, a proper privacy metric should protect every training instance against membership disclosure. However, this cannot be achieved by assessing the similarity between synthetic data and real data. To elucidate this point, consider a simple synthesizer that generates data by randomly adding a constant to the real attributes or aggregating some attributes’ values. In such a case, the similarity-based metrics may misleadingly indicate a low privacy risk since the synthetic data does not replicate any individual record. However, the synthesizer remains vulnerable since the membership disclosure risk of the aggregated records is alarmingly high.
Another concern of the similarity-based privacy metrics is the risk of likability. Since these metrics are computed by comparing the closeness with real training data, using these metrics in the synthesizing process (e.g., trying to decrease DCR during synthesis) provides another channel to leak information about the training data. For example, recent research [19] has demonstrated the feasibility of reconstructing significant portions of training data, when given black-box access to a data synthesizer and an oracle for computing the DCR of synthesized datasets. This revelation serves as a stark warning to practitioners who rely on such metrics.
Membership Inference Attacks. Another way to empirically verify the privacy of models is membership inference attacks [55], which has been extensively studied on discriminative models. However, membership inference is generally harder to mount on generative models than it is on discriminative ones since there are no clear confidence signals to detect overfitting and infer membership [25, 62]. Some studies [57] try to facilitate the problems by utilizing handcrafted features extracted from synthetic data distribution to train the shadow models and attack models. Another work [28] approximates the likelihood of target records with density estimation and uses it as the confidence score for attacks. However, as pointed out in [21] and [53], these attacks either depend on unrealistic settings or are too weak to differentiate the membership disclosure risks among various synthesizers [15]. Therefore, existing synthesis algorithms rarely use membership inference attacks for privacy evaluation [53].
4.2 Proposed Metric: Membership Disclosure Score
We propose a new privacy metric for data synthesis, termed membership disclosure score (MDS), to directly estimate the membership disclosure risks without relying on the effectiveness of attacks. We first quantify the disclosure risk of the target record $x$ as follows:
$\mathrm{DS}(x,\mathsf{A},\mathcal{D})\coloneqq\underset{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime}\subset\mathcal{D}\\ {x\in\mathcal{H},x\notin\mathcal{H}^{\prime}}\end{subarray}}{\mathbb{E}}\big{[}||\text{NN}(\mathcal{S},x)-\text{NN}(\mathcal{S}^{\prime},x)||_{1}\big{]},$ (6)
where $\mathcal{S}$ and $\mathcal{S}^{\prime}$ represent the synthetic datasets when the synthesizer $\mathsf{A}$ uses the subsets $\mathcal{H}$ and $\mathcal{H}^{\prime}$ respectively as the input to output a data generator. The term $\text{NN}(\mathcal{S},x)$ denotes the nearest neighbor (under $l_{1}$ distance) for record $x$ within the synthetic dataset $\mathcal{S}$. We also use $l_{1}$ distance [^6] Empirically we find that the difference between using $l_{1}$ and $l_{2}$ distance for computation is negligible. to measure the discrepancy resulting from training the model with or without the record $x$. Note that the disclosure score is calculated based on the difference between neighboring synthetic data, which ensures that real data information remains undisclosed. Accordingly, the membership disclosure risk of synthesizer $\mathsf{A}$ can be defined as:
Definition 2
(Membership Disclosure Score) Let $\mathrm{DS}(x,\mathsf{A},\mathcal{D})$ be the disclosure risk of record $x$ in real data $\mathcal{D}$ for synthesis algorithm $\mathsf{A}$. The membership disclosure score of the data synthesis algorithm $\mathsf{A}$ is given by:
$\mathrm{MDS}(\mathsf{A})\coloneqq\max_{x\in\mathcal{D}}~{}\mathrm{DS}(x,\mathsf{A},\mathcal{D}).$ (7)
Here we omit the notion of dataset $\mathcal{D}$ for simplicity. This metric signifies the maximum disclosure risk across the entire training dataset, which indicates that MDS is a worst-case privacy measurement. Intuitively, for a privacy-preserving synthesizer, the change of output (synthetic data) should remain insignificant by any individual training samples, resulting in a low MDS.
Directly computing this score entails retraining the model for each individual record, which is computationally prohibitive. Alternatively, we train $m$ models once on random subsets of $\mathcal{D}$, and compute the disclosure score for all instances using the same set of $m$ synthesis models. This approach provides an efficient and practical approach to estimating MDS in real-world scenarios.
Our definition of MDS aligns closely with concepts of memorization in neural networks [17, 67] and the leave-one-out notion of stability in machine learning [8]. However, it diverges in three crucial ways: (i) We measure the worst-case discourse risk as the privacy metric, whereas other studies focus on the difference of individuals or average cases. (ii) Our work specifically addresses privacy concerns in data synthesis, as opposed to other studies that explore discriminative models like classification. (iii) Our approach emphasizes the discrepancy caused by the presence or absence of target data in training, in contrast to other works that highlight performance gains from adding samples to the training set.
5 Utility Evaluation
In this section, we examine the existing utility measurements and propose two robust evaluation metrics for data synthesis.
5.1 Existing Metrics and Limitations
Machine Learning Efficacy. Machine learning efficacy has emerged as the predominant utility metric for data synthesis [32, 65, 68, 31]. Specifically, it first forms a classification (regression) problem by treating the target attribute as labels. Then, it computes the average performance of certain classifiers or regressors in a “train on synthetic data, test on real data” (TSTR) framework [30]. This metric, ideally, should reflect the synthetic data’s ability to preserve the real data distribution through test performance.
Shortfalls of Machine Learning Efficacy. The computation of this metric involves the selection of machine learning models for performance evaluation. It is known that tabular data performs differently when evaluated with various machine learning models [32]. Consequently, selecting different models for comparison can lead to divergent outcomes, complicating the process of fair assessment. Furthermore, it is expected that high-quality synthetic data should closely mimic the performance behavior of real data across a spectrum of machine learning models, which is pivotal for model selection. However, machine learning efficacy, as an average-based metric, fails to capture these performance discrepancies.
5.2 Proposed Metrics: MLA and Query Error
Machine Learning Affinity (MLA). The inconsistent performance of different machine learning models is caused by the distribution shift of synthetic data. It is known that machine learning models vary in their sensitivity to such shifts, leading to divergent levels of performance degradation [37]. To accurately reflect these performance fluctuations on synthetic data, we introduce an interpretable, easy-to-compute metric called machine learning affinity. It measures the relative accuracy difference between synthetic and real data under the TSTR framework. More formally we define:
Definition 3
[…]
Instead of directly averaging the performance on synthetic data, we measure the relative performance discrepancy over various machine learning models. Therefore, a low affinity not only indicates the effectiveness of the synthetic data for machine learning tasks but also represents similar performance behaviors with real ones.
Query Error. Range query and point query are fundamental tasks for tabular data analysis [10]. However, they are often overlooked in the evaluation of most state-of-the-art synthesizers [65, 68, 7, 60]. To accommodate both range query and point query, we follow [41] to define the query error based on the frequency of random $k$-way queries:
Definition 4
[…]
6 SynMeter
This section introduces SynMeter, a modular toolkit tailored for assessing data synthesis algorithms across three dimensions: fidelity, privacy, and utility.
6.1 Tuning Objective
Most synthesizers do not provide details about how to tune the hyperparameters [65, 60, 34, 41]. Instead, default settings are used for evaluations, which may not always yield optimal results, leading to subpar performance and skewed comparisons. To address this issue, we introduce the model tuning phase within our evaluation framework to facilitate the selection of hyperparameters for synthesizers. More precisely, we propose a novel and unified tuning objective, $\mathcal{L}$, to steer the tuning process:
$\mathcal{L}(\mathsf{A})=\alpha_{1}\text{Fidelity}(\mathsf{A})+\alpha_{2}\text{MLA}(\mathsf{A})+\alpha_{3}\text{QueryError}(\mathsf{A}).$
Here, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are coefficients assigned to fidelity, machine learning affinity, and query error. These coefficients enable the selection of optimal hyperparameters by assessing the fidelity and utility of synthetic data against the validation dataset $\mathcal{D}_{\text{val}}$. In our evaluation, we set all coefficients to 1, since we find that it can significantly improve the quality of the synthetic data compared to the default configurations. For instance, this approach can boost the fidelity of TabDDPM by 13% and its utility by at least 11% (refer to Section 7.6 for detailed comparisons). Practitioners can adjust these coefficients based on the specific requirements of their applications.
6.2 Implementation
SynMeter’s design is inherently modular, offering significant flexibility to incorporate new synthesizers and datasets. As depicted in Figure 1, SynMeter comprises four modules, each module is implemented with an abstract interface for any synthesis algorithm. We envisage that SynMeter can be used for the following purposes:
  • With eight state-of-the-art synthesis algorithms implemented, SynMeter facilitates data owners to tune, train, and select different synthesizers for data publishing.
  • It serves as a benchmark for data synthesis, providing principled evaluation metrics for comparative studies.
  • The modular architecture, connected via abstract interfaces, allows for easy integration of any synthesizers and enables adaptation into other domains (e.g., genomic data synthesis), by re-implementing module-specific processing functions.
Note that while several data synthesis libraries exist [53, 49], SynMeter is the only one that provides holistic data synthesis pipelines with principled metrics.
7 Experiments
Based on the proposed evaluation framework and metrics, in this section, we present a series of comprehensive experiments to answer the following question:
  • RQ1: How do the various synthesizers perform under the principled assessment? What are the new findings?
  • RQ2: Why do these methods work well (or not so well) on certain properties? How can our metrics help for in-depth analysis?
  • RQ3: How effective are our proposed metrics and evaluation framework?
Figure 1: Overview of SynMeter.
Figure 2: Average ranking comparison for six HP synthesizers. Each vertex is the average ranking of the method across 12 datasets (outer means higher rank), and each axis is the evaluation metric. “Fidelity($\mathcal{D}_{\text{train}}$)” denotes the fidelity is evaluated on training dataset, and “Fidelity($\mathcal{D}_{\text{test}}$)” means it tested on test dataset.
Figure 3: Visualization comparisons of synthetic data generated by six HP synthesizers on Bean dataset. The blue points are real data, and the orange ones are synthetic data.
7.1 Data Synthesis Models
We study a wide range of state-of-the-art data synthesizers, from statistical methods to deep generative models. We select them as they are either generally considered to perform best in practice [43, 44], widely used [65, 47], or recently emerged [32, 7, 60]. These synthesizers can be categorized into two groups: heuristic private (HP) synthesizers and DP synthesizers.
HP Synthesizers. Synthesizers in this category are developed without integrating DP:
  • CTGAN [65] is one of the most widely used HP synthesis algorithms. It utilizes generative adversarial networks to learn the tabular data distributions. Training techniques like conditional generation and Wasserstein loss [24] are used.
  • TVAE [65] is the state-of-the-art variational autoencoder for tabular data generation, which uses mode-specific normalization to overcome the non-Gaussian and multimodal problems of continuous distribution in tabular data.
  • TabDDPM [32] is the state-of-the-art diffusion model for data synthesis. It leverages the Gaussian diffusion process and the multinomial diffusion process to model continuous and discrete distributions respectively.
  • GReaT [7] utilizes the large language model (LLM) for data synthesis. It converts records to textual representations for LLM and generates synthetic data with prompts.
DP Synthesizers. These methods are either inherently designed with DP or are adaptations of HP models with additional mechanisms to offer provable privacy guarantees:
  • PGM [39] is the state-of-the-art DP synthesizer, which uses probabilistic graphical models [41] to learn the dependence of low-dimensional marginals for data synthesis. It won the NIST Differential Privacy Synthetic Data Challenge [43].
  • PrivSyn [70] is a non-parametric DP synthesizer, which iteratively updates the synthetic dataset to make it match the target noise marginals. This method also shows strong performance in NIST competitions [43, 44].
  • PATE-GAN [30] shares a similar architecture with CTGAN, but leverages the Private Aggregation of Teacher Ensembles (PATE) mechanism [47] to offer DP guarantees.
  • TableDiffusion [60] is a newly proposed diffusion model for tabular data synthesis, which uses Differentially Private Stochastic Gradient Descent (DP-SGD) to enforce privacy.
Note that all DP synthesizers can be adapted to the HP scenario by either using their HP equivalents (i.e., CTGAN as the HP model for PATE-GAN, TabDDPM for TableDiffusion) or setting the privacy budget to infinity (i.e., PGM and PrivSyn). However, some HP synthesizers, such as TVAE and GReaT, do not have corresponding DP variants. Thus, we only assess their performance within the context of HP models.
Our goal is not to benchmark all possible synthesizers but to focus on the best-known and broad spectrum of state-of-the-art synthesizers. Hence, some synthesizers [71, 40, 34, 31] are omitted as they are considered the variants of the algorithms we are assessing. Some approaches (e.g., TabDDPM, GReaT) are nearly simultaneously introduced, limiting opportunities for an extensive comparison. Our evaluation fills this gap by providing the first comprehensive evaluation of all state-of-the-art in a standardized setting.
Figure 4: Average ranking comparison for four DP synthesizers. Each vertex is the average ranking of the method across 12 datasets (outer means higher rank), and each axis is the evaluation metric. All methods offer provable privacy guarantees (thus all rank 1st).
Figure 5: Visualization comparisons of synthetic data generated by four DP synthesizers on Bean dataset.
Table 3: List of datasets used for evaluation.
Name # Train # Val # Test # Num # Cat Task type
Adult $20838$ $5210$ $6513$ $6$ $9$ Binclass
Shoppers $7891$ $1973$ $2466$ $10$ $8$ Binclass
Phishing $7075$ $1769$ $2211$ 00 $31$ Binclass
Magic $12172$ $3044$ $3804$ $10$ $1$ Binclass
Faults $1241$ $311$ $389$ $24$ $4$ Multiclass(7)
Bean $8710$ $2178$ $2723$ $16$ $1$ Multiclass(7)
Obesity $1350$ $338$ $423$ $8$ $9$ Multiclass(7)
Robot $3491$ $873$ $1092$ $24$ $1$ Multiclass(4)
Abalone $2672$ $668$ $836$ $8$ $1$ Regression
News $25372$ $6343$ $7929$ $46$ $14$ Regression
Insurance $856$ $214$ $268$ $3$ $4$ Regression
Wine $3134$ $784$ $980$ $12$ 00 Regression
7.2 Experimental Setups
Datasets. For systematic investigation of the performance of synthesis algorithms, we consider a diverse set of 12 real-world public datasets. These datasets have various sizes, nature, attributes, and distributions. We explicitly divide datasets into training and test with a ratio of 8:2, then split 20% of the training dataset as the validation set, which is used for model tuning. All datasets are publicly available and widely used [71, 32], and the full list is detailed in Table 3.
Evaluation Details. We use proposed fidelity, privacy, and utility metrics for evaluation. We first tune the synthesizers with our proposed tuning objective using training and validation data and train them with optimal performance. Then, synthetic data are generated by generator $\mathsf{G}$ for evaluation, where we test 10 times and report the mean and standard deviation as the final score. The implementation details of each metric are as follows:
  • We use all the one-way and two-way marginals for fidelity computation and use the mean as the final score. The Wasserstein distance for univariate distributions has a closed form and can be calculated easily. For two-way marginals, we use OPT package [18] to solve the optimal problem described in Equation 1. When the cost matrix is large ($n\geq 5,000$), we randomly sample half of the elements in the probability distributions for effective calculation.
  • To estimate the membership disclosure score for privacy evaluation, we train 80 synthesis models with randomly sampled data, and use them to compute the disclosure score for all training samples.
  • We utilize eight machine learning models for utility evaluations: SVM, Logistic Regression (or Ridge Regression), Decision Tree, Random Forest, Multilayer Perceptron (MLP), XGBoost [11], CatBoost [52], and Transformers [22]. Each model is extensively tuned on real training data to ensure optimal performance (see Appendix C for details). Performance on classification and regression is evaluated by the F1 score and RMSE, respectively.
  • We randomly construct 1,000 3-way query conditions and conduct range (point) queries for both synthetic and real data to compute the query errors.
7.3 Implementation Details
We use implementations from the original papers and make some small modifications to integrate them into our evaluation framework. For PGM and PrivSyn, we discretize numerical attributes into discrete ones by binning and setting the privacy budget to infinity when compared with HP synthesis algorithms. We use the same fixed random seeds for each model to improve reproducibility. To tune the hyperparameters of data synthesis algorithms and the machine learning models, we use the Optuna library [1]. The hyperparameter search space of each model is detailed in Appendix A.
7.4 Overall Evaluation (RQ1)
Overview. Figure 2 and Figure 4 report the overview ranking results for HP and DP synthesizers, respectively. We set the privacy budget $\epsilon$ to 1 for all DP synthesis algorithms. For HP synthesizers, the diffusion model (i.e., TabDDPM) exhibits superior fidelity and utility, albeit at the expense of compromising privacy. Statistical methods like PrivSyn also achieve impressive fidelity, surpassing complex deep generative models like TVAE and GReaT. Conversely, CTGAN, the most popular HP synthesizer, shows the least satisfactory results in synthetic data quality but offers better privacy protection.
In terms of DP synthesizers, statistical methods continue to perform well in terms of fidelity and utility. However, the performance of deep generative models drops significantly to ensure differential privacy. Even the strongest diffusion-based model (i.e., TableDiffusion) underperforms statistical approaches by a large margin, which starkly contrasts with its performance in the HP context, indicating a pronounced impact of privacy constraints on deep generative models.
We also visualize the synthetic and real data by using the t-SNE technique [63], as depicted in Figure 3and Figure 5. The visualizations generally reveal a considerable overlap between synthetic and real data points, demonstrating the synthesizers’ ability to generate data resembling the original. One interesting observation is that PrivSyn is the only method that does not have clear cluster patterns. This is because PrivSyn is the only non-parametric approach that directly modifies the synthetic records to match the marginal distribution, potentially resulting in less realistic records.
Table 4: Fidelity evaluation of data synthesis on train data $\mathcal{D}_{\text{train}}$. The privacy budget $\epsilon$ of HP synthesizers is set as $\infty$ (the top part), and the budget for the DP synthesizers is set as $1$ (the bottom part). The best result is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.186\scriptscriptstyle\pm\scriptstyle.010$ $0.092\scriptscriptstyle\pm\scriptstyle.002$ $0.019\scriptscriptstyle\pm\scriptstyle.001$ $0.037\scriptscriptstyle\pm\scriptstyle.002$ $0.056\scriptscriptstyle\pm\scriptstyle.002$ $0.040\scriptscriptstyle\pm\scriptstyle.002$ $0.041\scriptscriptstyle\pm\scriptstyle.001$ $0.050\scriptscriptstyle\pm\scriptstyle.002$ $0.037\scriptscriptstyle\pm\scriptstyle.002$ $0.060\scriptscriptstyle\pm\scriptstyle.001$ $0.038\scriptscriptstyle\pm\scriptstyle.005$ $0.066\scriptscriptstyle\pm\scriptstyle.001$
PrivSyn $0.024\scriptscriptstyle\pm\scriptstyle.001$ $0.030\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.010\scriptscriptstyle\pm\scriptstyle.001}\mathbf{}$ $0.015\scriptscriptstyle\pm\scriptstyle.003$ $0.064\scriptscriptstyle\pm\scriptstyle.006$ $0.035\scriptscriptstyle\pm\scriptstyle.002$ $0.034\scriptscriptstyle\pm\scriptstyle.002$ $0.065\scriptscriptstyle\pm\scriptstyle.012$ $0.024\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.018\scriptscriptstyle\pm\scriptstyle.001}$ $0.033\scriptscriptstyle\pm\scriptstyle.002$ $0.017\scriptscriptstyle\pm\scriptstyle.000$
TVAE $0.085\scriptscriptstyle\pm\scriptstyle.002$ $0.156\scriptscriptstyle\pm\scriptstyle.001$ $0.024\scriptscriptstyle\pm\scriptstyle.001$ $0.021\scriptscriptstyle\pm\scriptstyle.003$ $0.055\scriptscriptstyle\pm\scriptstyle.007$ $0.047\scriptscriptstyle\pm\scriptstyle.006$ $0.055\scriptscriptstyle\pm\scriptstyle.004$ $0.053\scriptscriptstyle\pm\scriptstyle.001$ $0.048\scriptscriptstyle\pm\scriptstyle.003$ $0.081\scriptscriptstyle\pm\scriptstyle.001$ $0.078\scriptscriptstyle\pm\scriptstyle.007$ $0.039\scriptscriptstyle\pm\scriptstyle.000$
CTGAN $0.059\scriptscriptstyle\pm\scriptstyle.001$ $0.062\scriptscriptstyle\pm\scriptstyle.001$ $0.062\scriptscriptstyle\pm\scriptstyle.002$ $0.157\scriptscriptstyle\pm\scriptstyle.006$ $0.133\scriptscriptstyle\pm\scriptstyle.004$ $0.139\scriptscriptstyle\pm\scriptstyle.005$ $0.072\scriptscriptstyle\pm\scriptstyle.002$ $0.106\scriptscriptstyle\pm\scriptstyle.003$ $0.049\scriptscriptstyle\pm\scriptstyle.004$ $0.040\scriptscriptstyle\pm\scriptstyle.001$ $0.090\scriptscriptstyle\pm\scriptstyle.004$ $0.033\scriptscriptstyle\pm\scriptstyle.001$
TabDDPM $\mathbf{0.020\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.022\scriptscriptstyle\pm\scriptstyle.001}$ $0.015\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.011\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.026\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.017\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.004}$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.028\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.011\scriptscriptstyle\pm\scriptstyle.000}$
[…]
PGM ($\epsilon=1$) $0.198\scriptscriptstyle\pm\scriptstyle.013$ $0.103\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.023\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.042\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.086\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.048\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.063\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.065\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.052\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.062\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.071\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.068\scriptscriptstyle\pm\scriptstyle.001}$
PrivSyn ($\epsilon=1$) $\mathbf{0.045\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.077\scriptscriptstyle\pm\scriptstyle.005}$ $0.033\scriptscriptstyle\pm\scriptstyle.002$ $0.052\scriptscriptstyle\pm\scriptstyle.003$ $0.228\scriptscriptstyle\pm\scriptstyle.007$ $0.142\scriptscriptstyle\pm\scriptstyle.007$ $0.167\scriptscriptstyle\pm\scriptstyle.009$ $0.169\scriptscriptstyle\pm\scriptstyle.021$ $0.127\scriptscriptstyle\pm\scriptstyle.009$ $0.070\scriptscriptstyle\pm\scriptstyle.002$ $0.124\scriptscriptstyle\pm\scriptstyle.006$ $0.156\scriptscriptstyle\pm\scriptstyle.004$
PATE-GAN ($\epsilon=1$) $0.139\scriptscriptstyle\pm\scriptstyle.001$ $0.176\scriptscriptstyle\pm\scriptstyle.002$ $0.173\scriptscriptstyle\pm\scriptstyle.002$ $0.153\scriptscriptstyle\pm\scriptstyle.005$ $0.204\scriptscriptstyle\pm\scriptstyle.003$ $0.520\scriptscriptstyle\pm\scriptstyle.006$ $0.086\scriptscriptstyle\pm\scriptstyle.003$ $0.477\scriptscriptstyle\pm\scriptstyle.002$ $0.331\scriptscriptstyle\pm\scriptstyle.005$ $0.065\scriptscriptstyle\pm\scriptstyle.002$ $0.385\scriptscriptstyle\pm\scriptstyle.003$ $0.251\scriptscriptstyle\pm\scriptstyle.000$
TableDiffusion ($\epsilon=1$) $0.180\scriptscriptstyle\pm\scriptstyle.002$ $0.209\scriptscriptstyle\pm\scriptstyle.002$ $0.123\scriptscriptstyle\pm\scriptstyle.002$ $0.132\scriptscriptstyle\pm\scriptstyle.003$ $0.369\scriptscriptstyle\pm\scriptstyle.002$ $0.148\scriptscriptstyle\pm\scriptstyle.005$ $0.347\scriptscriptstyle\pm\scriptstyle.003$ $0.203\scriptscriptstyle\pm\scriptstyle.001$ $0.232\scriptscriptstyle\pm\scriptstyle.005$ $0.135\scriptscriptstyle\pm\scriptstyle.001$ $0.343\scriptscriptstyle\pm\scriptstyle.002$ $0.108\scriptscriptstyle\pm\scriptstyle.001$
  • 1
    GReaT cannot be applied to the News dataset because of the maximum length limit of large language models.
Table 5: Privacy evaluation (membership disclosure score) of HP synthesizers ($\epsilon=\infty$). The lower score means the synthesizer can provide better empirical membership privacy protection. The best result is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.131\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.133\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.108\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.028\scriptscriptstyle\pm\scriptstyle.001}$ $0.125\scriptscriptstyle\pm\scriptstyle.002$ $0.032\scriptscriptstyle\pm\scriptstyle.003$ $0.138\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.031\scriptscriptstyle\pm\scriptstyle.002}$ $0.113\scriptscriptstyle\pm\scriptstyle.003$ $0.041\scriptscriptstyle\pm\scriptstyle.001$ $0.242\scriptscriptstyle\pm\scriptstyle.002$ $0.120\scriptscriptstyle\pm\scriptstyle.003$
PrivSyn $0.187\scriptscriptstyle\pm\scriptstyle.002$ $0.164\scriptscriptstyle\pm\scriptstyle.001$ $0.139\scriptscriptstyle\pm\scriptstyle.003$ $0.033\scriptscriptstyle\pm\scriptstyle.002$ $0.083\scriptscriptstyle\pm\scriptstyle.001$ $0.037\scriptscriptstyle\pm\scriptstyle.003$ $0.132\scriptscriptstyle\pm\scriptstyle.002$ $0.078\scriptscriptstyle\pm\scriptstyle.001$ $0.113\scriptscriptstyle\pm\scriptstyle.003$ $0.059\scriptscriptstyle\pm\scriptstyle.002$ $0.240\scriptscriptstyle\pm\scriptstyle.001$ $0.071\scriptscriptstyle\pm\scriptstyle.003$
TVAE $0.199\scriptscriptstyle\pm\scriptstyle.003$ $0.186\scriptscriptstyle\pm\scriptstyle.002$ $0.172\scriptscriptstyle\pm\scriptstyle.001$ $0.030\scriptscriptstyle\pm\scriptstyle.005$ $0.093\scriptscriptstyle\pm\scriptstyle.002$ $0.042\scriptscriptstyle\pm\scriptstyle.001$ $0.155\scriptscriptstyle\pm\scriptstyle.003$ $0.082\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.045\scriptscriptstyle\pm\scriptstyle.001}$ $0.055\scriptscriptstyle\pm\scriptstyle.003$ $0.217\scriptscriptstyle\pm\scriptstyle.002$ $0.046\scriptscriptstyle\pm\scriptstyle.001$
CTGAN $\mathbf{0.072\scriptscriptstyle\pm\scriptstyle.002}$ $0.171\scriptscriptstyle\pm\scriptstyle.003$ $0.149\scriptscriptstyle\pm\scriptstyle.001$ $030$003 $\mathbf{0.075\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.031\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.112\scriptscriptstyle\pm\scriptstyle.001}$ $0.052\scriptscriptstyle\pm\scriptstyle.003$ $0.045\scriptscriptstyle\pm\scriptstyle.002$ $0.051\scriptscriptstyle\pm\scriptstyle.005$ $\mathbf{0.176\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.024\scriptscriptstyle\pm\scriptstyle.001}$
TabDDPM $0.550\scriptscriptstyle\pm\scriptstyle.001$ $0.230\scriptscriptstyle\pm\scriptstyle.002$ $0.720\scriptscriptstyle\pm\scriptstyle.003$ $0.031\scriptscriptstyle\pm\scriptstyle.001$ $0.130\scriptscriptstyle\pm\scriptstyle.002$ $0.064\scriptscriptstyle\pm\scriptstyle.003$ $0.543\scriptscriptstyle\pm\scriptstyle.001$ $0.164\scriptscriptstyle\pm\scriptstyle.002$ $0.141\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.030\scriptscriptstyle\pm\scriptstyle.001}$ $0.383\scriptscriptstyle\pm\scriptstyle.002$ $0.120\scriptscriptstyle\pm\scriptstyle.003$
GReaT $0.495\scriptscriptstyle\pm\scriptstyle.002$ $0.236\scriptscriptstyle\pm\scriptstyle.003$ $0.568\scriptscriptstyle\pm\scriptstyle.001$ $0.082\scriptscriptstyle\pm\scriptstyle.002$ $0.125\scriptscriptstyle\pm\scriptstyle.003$ $0.036\scriptscriptstyle\pm\scriptstyle.004$ $0.357\scriptscriptstyle\pm\scriptstyle.002$ $0.181\scriptscriptstyle\pm\scriptstyle.003$ $0.065\scriptscriptstyle\pm\scriptstyle.001$ $0.536\scriptscriptstyle\pm\scriptstyle.002$ $0.072\scriptscriptstyle\pm\scriptstyle.003$
Fidelity Evaluation. We assess the data synthesis fidelity by applying the Wasserstein distance to both training $\mathcal{D}_{\text{train}}$ and test data $\mathcal{D}_{\text{test}}$. The results in Table 4 and Appendix Table D17 demonstrate that TabDDPM leads among HP synthesizers, while statistical methods (e.g., PGM) excel within the DP synthesizers category. To achieve differential privacy, all deep generative models experience a significant fidelity loss, whereas statistical methods maintain stable performance. This stability is further highlighted when comparing performance across training and test data, where statistical methods exhibit a lesser performance drop compared to deep generative models. This suggests that statistical methods are more robust in data synthesis. A closer examination of the fidelity of each type of attribute (detailed in Appendix Section D.1) reveals that statistical methods generally achieve better fidelity performance on categorical attributes but fall short with numerical ones. This discrepancy can be attributed to the inherent design of statistical methods for discrete spaces, making them less effective for numerical data. Despite DDPM being enhanced with the multinomial diffusion process to better model discrete distributions, it still does not match the performance of statistical methods in terms of categorical fidelity.
Privacy Evaluation. Table 5 presents the privacy assessment for HP synthesizers. Contrary to the fidelity evaluation findings, CTGAN, which exhibited the lowest fidelity performance, provides superior privacy protections against membership disclosure. Statistical methods, such as PGM, also demonstrate commendable empirical privacy protections. Nonetheless, the unsatisfied results of strong synthesis algorithms like TabDDPM and GReaT reveal their vulnerability to membership disclosure.
Utility Evaluation. The utility of data synthesis is assessed by performing downstream tasks on the synthetic datasets and measuring with the proposed metrics (machine learning affinity and query errors), as shown in Table D28 and Table D29 (in Appendix). In the context of machine learning tasks, TabDDPM excels among HP synthesizers, which contributes to its class-conditional framework that learns label dependencies during its training process. However, this advantage diminishes when adding random noise to ensure privacy, where PGM takes the lead with its stable and superior performance. The outcomes for range (point) query tasks echo the fidelity evaluation results, where TabDDPM leads in scenarios devoid of provable privacy guarantees, and statistical methods can surpass other methods under DP constraints.
Figure 6: Impact of privacy budget $\epsilon$ on Bean dataset. “Train Fidelity” means the fidelity is evaluated on train dataset $\mathcal{D}_{\text{train}}$, and “Test Fidelity” denotes evaluated on test dataset $\mathcal{D}_{\text{test}}$. “MLA” is the proposed machine learning affinity.
7.5 In-depth Analysis (RQ2)
In this section, we delve into the underlying reasons for the observed performances. Specifically, we employ the proposed fidelity metrics as tools to auditing the learning process of different synthesizers and explore the impact of differing privacy budgets on different DP synthesizers.
(a) CTGAN
(b) TabDDPM
Figure 7: Auditing the learning process of CTGAN and TabDDPM with fidelity metrics on the Bean dataset. The fidelity of numerical-numerical marginals is omitted since the Bean dataset only has one numerical attribute.
Why Dose CTGAN Perform Poorly? CTGAN is widely recognized as a strong synthesis algorithm in many studies [32, 7, 71]. Despite its popularity, our evaluation reveals that it produces the lowest quality synthetic data. This discrepancy prompts an investigation into the reasons behind CTGAN’s apparent underachievement.
To address this, we scrutinize CTGAN’s learning trajectory, particularly evaluating the fidelity across different marginal types (categorical, numerical, and mixed) throughout training. Our findings illustrated in Figure 7(a), reveal an unexpected stagnation in the improvement of Wasserstein distances for both numerical and categorical marginals throughout the entire training period. As such, the quality of synthetic data completely relies on data preprocessing, which employs a variational Gaussian mixture model to approximate numerical data and conditional sampling for categorical attributes. The closer the data distribution to Gaussianity, the better the fidelity. Yet, most tabular data display complexities beyond simple Gaussian distributions [22], leading to CTGAN’s subpar performance. Furthermore, this phenomenon sheds light on CTGAN’s strong empirical privacy protections. The nearly random (and unlearnable) nature of the output does not depend on any specific training data presence. This aspect, while beneficial for privacy, underscores the challenges in generating high-fidelity synthetic data using CTGAN.
Why Dose TabDDPM Excel? A notable finding from our evaluations is TabDDPM’s capability to maintain high fidelity of low-dimensional marginals. This contradicts with previous statements that deep generative models generally fail for tabular data synthesis [40, 58]. To better illustrate this, we utilized fidelity metrics to audit TabDDPM’s learning process, as depicted in Figure 7(b). A rapid decrease of the Wasserstein distance for all types of marginals is observed during the training, indicating the ability of TabDDPM to model both numerical and categorical attributes. We contribute the efficacy of TabDDPM to the model architecture and categorical distribution modeling: Diffusion models are theoretically proved to minimize the Wasserstein distance between synthetic and real datasets [33]. This represents a methodological superiority compared to other generative models, which primarily aim at reducing the Kullback–Leibler divergence. Moreover, the integration of the multinomial diffusion process addresses the challenge of modeling discrete spaces for conventional diffusion models. However, TabDDPM faces serious privacy disclosure risks which are ignored by previous work [32], and directly applying differential privacy (i.e., TableDiffusion) would greatly impact the model’s ability to generate authentic data. Despite these challenges, we deem that diffusion-based methods represent a promising frontier for data synthesis.
Large Language Models Are Semantic-aware Synthesizers. We also notice that the recently emerged LLM-based synthesizer (i.e., GReaT) also shows competitive performance for data synthesis, which is capable of modeling semantic knowledge within attributes. This capability enables GReaT to excel in datasets characterized by explicit semantics. A good example is the Adult dataset, which comprises detailed personal information (e.g., age and income), and GReaT achieves the best performance on machine learning predictions among all HP synthesizers. Nevertheless, this advantage diminishes in cases where the attributes’ semantics are unclear. This is evident in the Phishing dataset, where all attributes’ values are preprocessed into integers and GReaT’s performance is notably weaker. Given the rapid development of LLM and the inherent rich semantics of most tabular data, it may become a new paradigm for realistic data synthesis.
The Impact of Privacy Budget. We conduct experiments on different privacy budgets to further analyze the impact of differential privacy for data synthesis. Specifically, we vary $\epsilon$ from $0.5$ to $16$, and evaluate the fidelity and utility of synthetic data, as shown in Figure 6. It was observed that statistical methods (e.g., PGM) demonstrate robust performance even with the small privacy budget ($\epsilon=0.5$), whereas deep generative models generally require a much larger privacy budget (e.g., 8) to yield similar outcomes. This is reasonable since statistical methods only rely on the estimation of a small set of marginals. In contrast, deep generative models aim to directly capture the entire joint distribution, making them inherently more susceptible to random perturbations.
Table 6: Average performance improvements (%) when tuning with our framework.
Models Fidelity $\uparrow$ Utility normal-$\uparrow$
$\mathcal{D}_{\text{Train}}$ $\mathcal{D}_{\text{Test}}$ MLA Query Error
PGM 0.33 0.34 17.35 3.39
PrivSyn 1.60 2.92 12.08 1.12
TVAE 1.06 0.67 5.29 2.67
CTGAN 9.87 9.60 0.57 8.63
PATE-GAN 6.27 8.48 0.75 7.04
TabDDPM 13.62 13.65 13.67 11.95
TableDiffusion 11.34 10.95 8.32 7.86
GReaT 3.84 9.21 1.14 1.77
7.6 Effectiveness of proposed Metrics (RQ3)
In this section, we investigate the efficacy of proposed metrics by conducting a series of comparison studies.
Impact of Model Tuning Phase. We initiated our investigation by examining the impact of the proposed tuning objective in Section 6.1. The data presented in Table 6 highlights the average performance improvements across all datasets. Generally, the tuning objective can consistently improve both the fidelity and utility of synthetic data for all synthesis algorithms. We also notice that the improvement in utility surpasses those in fidelity for statistical methods, while deep generative models exhibit the reverse trend. In addition, the tuning phrase is exceptionally beneficial for TabDDPM, yielding significant improvements in both fidelity and utility. Conversely, GReaT and TVAE show a greater resilience to model tuning.
Analysis of Membership Disclosure Score. We investigate the proposed membership disclosure score (MDS) by employing it to evaluate the empirical privacy protections of DP synthesizers. Specifically, we set the privacy budget $\epsilon$ from 1 to infinity and calculate the corresponding MDS. The results on the Adult and Phishing dataset are shown in Figure 8. It is observed that MDS can indicate the change in privacy budget, and a larger privacy budget would lead to higher membership disclosure risks (thus larger MDS). Interestingly, we find that statistical methods can provide reasonable empirical privacy protections, even without DP guarantees (i.e., $\epsilon=\infty$). On the contrary, deep generative models like TableDiffusion may need a smaller privacy budget to reduce empirical privacy risks. This may be because statistical methods are inherently more private since they only rely on statistics for synthesis. We leave it as an open question for future research.
8 Key Takeaways
Data synthesis has been advocated as an important approach for utilizing data while protecting data privacy. Despite the plethora of synthesis algorithms available, a comprehensive understanding of their strengths and weaknesses remains elusive due to the lack of principled evaluation metrics. In this paper, we examine and critique existing evaluation metrics, and introduce a principled and systematic framework for assessing data synthesis, as well as a new suite of evaluation criteria in terms of fidelity, privacy, and utility. We also provide a unified tuning objective that can consistently improve the performance of synthesis models and ensure that evaluations are grounded.
In the following, we summarize important insights obtained from our evaluations and provide some actionable advice to future data synthesis practitioners:
(a) Adult Dataset
(b) Phishing Dataset
Figure 8: Empirical privacy evaluation with membership disclosure score (MDS) on DP synthesizers.
  • No Free lunch in data synthesis. Currently, the performance gap between HP and DP synthesizers is quite large. No synthesizer can achieve the best performance while maintaining good privacy protections (even empirically).
  • Model tuning is indispensable for synthesizers. we find that the performance of all synthesizers can be boosted with our proposed tuning objective compared to the default settings. Therefore, tuning is crucial to reach the right conclusions.
  • Diffusion models are strong tabular synthesizers. We observe that diffusion models (i.e., TabDDPM) are capable of synthesizing highly authentic tabular data, significantly outperforming other synthesis algorithms. However, the privacy issue of diffusion models needs further investigation.
  • Statistical methods are robust (private) synthesizers. While many studies consider statistical methods to be weak to compare [68, 32, 34, 31], we observe that they achieve good and stable performance under DP constraints. We also find that statistical methods may be inherently more private than deep generative methods, which leaves an open question for further exploration.
  • Large language models are promising synthesizers. GReaT is the first LLM-based tabular synthesis, which utilizes fine-tuning of a pre-trained LLM on textually encoded tabular data. Despite the simplicity, it shows impressive performance on datasets that consist of rich semantic attributes.
  • CTGAN is too weak to be a practical synthesizer. Our analysis highlights that CTGAN falls short as an effective synthesizer, with no discernible improvement in synthetic data quality observed during its training phase. Therefore, a synthesizer’s performance should not be judged solely in comparison to CTGAN.
Our principled assessment leads us to conclude that recently emerged generative models achieve impressive performance for tabular data synthesis and offer new directions in this field. At the same time, several critical challenges are also revealed (e.g., privacy issues of diffusion models, gaps between DP and HP synthesizers, etc.). Our evaluation metrics and framework play a crucial role in elucidating the advancements in data synthesis as well as serve as a benchmark for future works.
Appendix
Appendix A Hyperparameters Search Spaces of synthesis Algorithms
In this paper, we tune and use the following synthesizers:
We set the privacy budget to infinity ($\epsilon=100,000,000$) when using DP synthesizers for HP evaluations.
Table A1: PGM [41] hyperparameters search space.
Parameter Distribution
Number of two-way marginals $\mathrm{Int}[10,50]$
Number of three-way marginals $\mathrm{Int}[5,20]$
Number of bins $\mathrm{Int}[5,20]$
Maximum number of iterations $\mathrm{Int}[3000,5000]$
Number of tuning trials 50
Table A2: PrivSyn [70] hyperparameters search space.
Parameter Distribution
Number of bins $\mathrm{Int}[5,20]$
Maximum number of iterations $\mathrm{Int}[10,100]$
Number of tuning trials 50
Table A3: TVAE [65] hyperparameters search space.
Parameter Distribution
Number of epochs $\mathrm{Int}[100,500]$
Batch size $\mathrm{Int}[500,5000]$
Loss factor $\mathrm{Float}[1,5]$
Embedding dimension $\mathrm{Int}[128,512]$
Compression dimension $\mathrm{Int}[128,512]$
Decompression dimension $\mathrm{Int}[128,512]$
$L_{2}$ regularization $\mathrm{LogUniform}[1e\text{-}6,1e\text{-}3]$
Number of tuning trials 50
Table A4: CTGAN [65] hyperparameters search space.
Parameter Distribution
Number of epochs $\mathrm{Int}[100,500]$
Batch size $\mathrm{Int}[500,5000]$
Embedding dimension $\mathrm{Int}[128,512]$
Generator dimension $\mathrm{Int}[128,512]$
Discriminator dimension $\mathrm{Int}[128,512]$
Learning rate of generator $\mathrm{LogUniform}[1e\text{-}5,1e\text{-}3]$
Learning rate of discriminator $\mathrm{LogUniform}[1e\text{-}5,1e\text{-}3]$
Number of tuning trials 50
Table A5: TabDDPM [32] hyperparameters search space.
Parameter Distribution
Number of layers $\mathrm{Int}[2,8]$
Embedding dimension $\mathrm{Int}[128,512]$
Number of diffusion timesteps $\mathrm{Int}[100,10000]$
Number of training iterations $\mathrm{Int}[5000,30000]$
Learning rate of discriminator $\mathrm{LogUniform}[1e\text{-}5,3e\text{-}3]$
Number of tuning trials 50
Table A6: GReaT [7] hyperparameters search space.
Parameter Distribution
Temperature $\mathrm{Float}[0.6,0.9]$
Number of fine-tuning epochs $\mathrm{Int}[100,300]$
Number of training iterations $\mathrm{Int}[5000,30000]$
Batch size $\mathrm{Int}[8,32]$
Number of tuning trials 20
Table A7: PATE-GAN [30] hyperparameters search space.
Parameter Distribution
Number of teachers $\mathrm{Int}[5,20]$
Number of generator layers $\mathrm{Int}[1,3]$
Number of discriminator layers $\mathrm{Int}[1,3]$
Generator dimension $\mathrm{Int}[50,200]$
Discriminator dimension $\mathrm{Int}[50,200]$
Number of iterations $\mathrm{Int}[1000,5000]$
Learning rate $\mathrm{LogUniform}[1e\text{-}5,1e\text{-}3]$
Number of tuning trials 50
Table A8: TableDiffusion [60] hyperparameters search space.
Parameter Distribution
Number of layers $\mathrm{Int}[1,6]$
Number of diffusion timesteps $\mathrm{Int}[3,20]$
Number of epochs $\mathrm{Int}[5,20]$
Batch size $\mathrm{Int}[128,1024]$
Noise prediction {True, False}
Learning rate $\mathrm{LogUniform}[1e\text{-}4,1e\text{-}2]$
Number of tuning trials 50
Appendix B Datasets
We use 12 real-world datasets for evaluations, varying in size, nature, and distribution. These datasets are all publicly available:
Appendix C Hyperparameters Search Spaces of Machine Learning Models
We tune and use the following machine learning models to evaluate the utility of synthesizers:
Table C9: SVM hyperparameters search space.
Parameter Distribution
Regularization parameter $\mathrm{LogUniform}[1e\text{-}5,1e\text{-}1]$
Kernel {linear, poly, rbf, sigmoid}
Number of tuning trials 50
Table C10: Logistic regression (Ridge regression) hyperparameters search space.
Parameter Distribution
Regularization $\mathrm{Float}[0,10]$
Maximum number of iterations $\mathrm{Int}[100,1000]$
Fit intercept {True, False}
Number of tuning trials 50
Table C11: Decision tree hyperparameters search space.
Parameter Distribution
Maximum depth $\mathrm{Int}[4,64]$
Minimum samples split $\mathrm{Int}[2,8]$
Minimum samples leaf $\mathrm{Int}[1,8]$
Number of tuning trials 50
Table C12: Random forest hyperparameters search space.
Parameter Distribution
Number of estimators $\mathrm{Int}[10,200]$
Maximum depth $\mathrm{Int}[4,64]$
Minimum samples split $\mathrm{Int}[2,8]$
Minimum samples leaf $\mathrm{Int}[1,8]$
Number of tuning trials 50
Table C13: MLP hyperparameters search space.
Parameter Distribution
Maximum number of iterations $\mathrm{Float}[50,200]$
Regularization $\mathrm{LogUniform}[1e\text{-}5,1e\text{-}1]$
Number of tuning trials 50
Table C14: XGBoost hyperparameters search space.
Parameter Distribution
Step size shrinkage $\mathrm{Float}[0.01,0.2]$
Minimum sum of instance weight $\mathrm{Int}[1,10]$
Maximum depth $\mathrm{Int}[3,20]$
Minimum loss reduction $\mathrm{Float}[0,1]$
Number of tuning trials 50
Table C15: CatBoost hyperparameters search space from [22].
Parameter Distribution
Max depth $\mathrm{UniformInt[3,10]}$
Learning rate $\mathrm{LogUniform}[1e\text{-}5,1]$
Bagging temperature $\mathrm{Uniform}[0,1]$
$L_{2}$ leaf reg $\mathrm{LogUniform}[1,10]$
Leaf estimation iterations $\mathrm{UniformInt}[1,10]$
Number of tuning trials 50
Table C16: Transformer hyperparameters space from [22].
Parameter Distribution
Number of layers $\mathrm{UniformInt}[1,8]$
Layer size $\mathrm{Int}\{64,128,256,512,1024\}$
Dropout $\{0,\mathrm{Uniform}[0,0.5]\}$
Learning rate $\mathrm{LogUniform}[1e\text{-}5,1e\text{-}2]$
Weight decay $\{0,\mathrm{LogUniform}[1e\text{-}6,1e\text{-}3]\}$
Number of tuning trials 50
Appendix D Complete Experimental Results
D.1 Complete Fidelity Results
Here we include the remaining fidelity results in our evaluation. The overall fidelity evaluation on test data is demonstrated in Table D17.
In addition, the complete fidelity results consist of the Wasserstein distance for two one-way marginals (categorical and numerical distributions) and three two-way marginals (categorical-categorical distribution, numerical-numerical distribution, categorical-numerical distribution).
For fidelity on test data $\mathcal{D}_{\text{train}}$, the detailed Wasserstein distances for one-way marginals are shown in Table D18 and Table D19. The Wasserstein distances for two-way marginals are shown in Table D20, Table D21 and Table D22.
For fidelity on test data $\mathcal{D}_{\text{test}}$, the detailed Wasserstein distances for one-way marginals are shown in Table D23 and Table D24. The Wasserstein distances for two-way marginals are shown in Table D25, Table D26 and Table D27.
Table D17: Fidelity evaluation of data synthesis on test data $\mathcal{D}_{\text{test}}$. The privacy budget $\epsilon$ of HP synthesis algorithms is set as $\infty$ (the top part), and the budget for DP synthesis algorithm is set as $1$ (the bottom part). The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.172\scriptscriptstyle\pm\scriptstyle.004$ $0.098\scriptscriptstyle\pm\scriptstyle.002$ $0.026\scriptscriptstyle\pm\scriptstyle.001$ $0.039\scriptscriptstyle\pm\scriptstyle.002$ $0.089\scriptscriptstyle\pm\scriptstyle.006$ $0.044\scriptscriptstyle\pm\scriptstyle.003$ $0.062\scriptscriptstyle\pm\scriptstyle.003$ $0.055\scriptscriptstyle\pm\scriptstyle.003$ $0.062\scriptscriptstyle\pm\scriptstyle.008$ $0.050\scriptscriptstyle\pm\scriptstyle.004$ $0.083\scriptscriptstyle\pm\scriptstyle.009$ $0.075\scriptscriptstyle\pm\scriptstyle.002$
PrivSyn $0.025\scriptscriptstyle\pm\scriptstyle.001$ $0.041\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.017\scriptscriptstyle\pm\scriptstyle.002}$ $0.015\scriptscriptstyle\pm\scriptstyle.002$ $0.079\scriptscriptstyle\pm\scriptstyle.007$ $0.037\scriptscriptstyle\pm\scriptstyle.003$ $0.053\scriptscriptstyle\pm\scriptstyle.005$ $0.054\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.032\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.018\scriptscriptstyle\pm\scriptstyle.001}$ $0.074\scriptscriptstyle\pm\scriptstyle.006$ $0.022\scriptscriptstyle\pm\scriptstyle.001$
TVAE $0.086\scriptscriptstyle\pm\scriptstyle.002$ $0.154\scriptscriptstyle\pm\scriptstyle.002$ $0.028\scriptscriptstyle\pm\scriptstyle.002$ $0.020\scriptscriptstyle\pm\scriptstyle.003$ $0.081\scriptscriptstyle\pm\scriptstyle.016$ $0.050\scriptscriptstyle\pm\scriptstyle.004$ $0.059\scriptscriptstyle\pm\scriptstyle.003$ $0.059\scriptscriptstyle\pm\scriptstyle.007$ $0.046\scriptscriptstyle\pm\scriptstyle.005$ $0.079\scriptscriptstyle\pm\scriptstyle.001$ $0.118\scriptscriptstyle\pm\scriptstyle.009$ $0.045\scriptscriptstyle\pm\scriptstyle.001$
CTGAN $0.061\scriptscriptstyle\pm\scriptstyle.003$ $0.061\scriptscriptstyle\pm\scriptstyle.002$ $0.069\scriptscriptstyle\pm\scriptstyle.001$ $0.150\scriptscriptstyle\pm\scriptstyle.004$ $0.133\scriptscriptstyle\pm\scriptstyle.007$ $0.139\scriptscriptstyle\pm\scriptstyle.005$ $0.085\scriptscriptstyle\pm\scriptstyle.004$ $0.109\scriptscriptstyle\pm\scriptstyle.009$ $0.066\scriptscriptstyle\pm\scriptstyle.005$ $0.040\scriptscriptstyle\pm\scriptstyle.001$ $0.116\scriptscriptstyle\pm\scriptstyle.008$ $0.034\scriptscriptstyle\pm\scriptstyle.001$
TabDDPM $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.031\scriptscriptstyle\pm\scriptstyle.001}$ $0.019\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.012\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.058\scriptscriptstyle\pm\scriptstyle.008}$ $\mathbf{0.016\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.043\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.028\scriptscriptstyle\pm\scriptstyle.004}$ $0.034\scriptscriptstyle\pm\scriptstyle.010$ $0.032\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.070\scriptscriptstyle\pm\scriptstyle.009}$ $\mathbf{0.017\scriptscriptstyle\pm\scriptstyle.001}$
GReaT $0.052\scriptscriptstyle\pm\scriptstyle.002$ $0.056\scriptscriptstyle\pm\scriptstyle.004$ $0.072\scriptscriptstyle\pm\scriptstyle.002$ $0.039\scriptscriptstyle\pm\scriptstyle.003$ $0.063\scriptscriptstyle\pm\scriptstyle.007$ $0.021\scriptscriptstyle\pm\scriptstyle.004$ $0.062\scriptscriptstyle\pm\scriptstyle.008$ $0.058\scriptscriptstyle\pm\scriptstyle.006$ $0.037\scriptscriptstyle\pm\scriptstyle.004$ $0.107\scriptscriptstyle\pm\scriptstyle.010$ $0.024\scriptscriptstyle\pm\scriptstyle.001$
PGM ($\epsilon=1$) $0.179\scriptscriptstyle\pm\scriptstyle.004$ $0.103\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.028\scriptscriptstyle\pm\scriptstyle.001}$ $0.042\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.112\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.048\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.075\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.072\scriptscriptstyle\pm\scriptstyle.007}$ $\mathbf{0.080\scriptscriptstyle\pm\scriptstyle.010}$ $0.051\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.093\scriptscriptstyle\pm\scriptstyle.006}$ $\mathbf{0.075\scriptscriptstyle\pm\scriptstyle.001}$
PrivSyn ($\epsilon=1$) $\mathbf{0.049\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.084\scriptscriptstyle\pm\scriptstyle.002}$ $030$003 $\mathbf{0.031\scriptscriptstyle\pm\scriptstyle.003}$ $0.236\scriptscriptstyle\pm\scriptstyle.017$ $0.128\scriptscriptstyle\pm\scriptstyle.010$ $0.154\scriptscriptstyle\pm\scriptstyle.013$ $0.177\scriptscriptstyle\pm\scriptstyle.011$ 0.111±.011plus-or-minus0.111.0110.$1$011 $\mathbf{0.044\scriptscriptstyle\pm\scriptstyle.001}$ $0.152\scriptscriptstyle\pm\scriptstyle.011$ $0.130\scriptscriptstyle\pm\scriptstyle.005$
PATE-GAN ($\epsilon=1$) $0.139\scriptscriptstyle\pm\scriptstyle.002$ $0.171\scriptscriptstyle\pm\scriptstyle.002$ $0.173\scriptscriptstyle\pm\scriptstyle.002$ $0.155\scriptscriptstyle\pm\scriptstyle.005$ $0.215\scriptscriptstyle\pm\scriptstyle.004$ $0.523\scriptscriptstyle\pm\scriptstyle.004$ $0.089\scriptscriptstyle\pm\scriptstyle.004$ $0.478\scriptscriptstyle\pm\scriptstyle.007$ $0.353\scriptscriptstyle\pm\scriptstyle.009$ $0.061\scriptscriptstyle\pm\scriptstyle.002$ $0.386\scriptscriptstyle\pm\scriptstyle.011$ $0.250\scriptscriptstyle\pm\scriptstyle.003$
TableDiffusion ($\epsilon=1$) $0.179\scriptscriptstyle\pm\scriptstyle.002$ $0.210\scriptscriptstyle\pm\scriptstyle.002$ $0.121\scriptscriptstyle\pm\scriptstyle.002$ $0.132\scriptscriptstyle\pm\scriptstyle.005$ $0.390\scriptscriptstyle\pm\scriptstyle.004$ $0.149\scriptscriptstyle\pm\scriptstyle.004$ $0.338\scriptscriptstyle\pm\scriptstyle.005$ $0.203\scriptscriptstyle\pm\scriptstyle.002$ $0.226\scriptscriptstyle\pm\scriptstyle.007$ $0.128\scriptscriptstyle\pm\scriptstyle.001$ $0.366\scriptscriptstyle\pm\scriptstyle.008$ $0.098\scriptscriptstyle\pm\scriptstyle.001$
Table D18: Fidelity evaluation of one-way categorical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.015\scriptscriptstyle\pm\scriptstyle.002$ $0.012\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.005\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.004\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.013\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.011\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.007\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.008\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.005\scriptscriptstyle\pm\scriptstyle.000}$ $0.018\scriptscriptstyle\pm\scriptstyle.007$
PrivSyn $\mathbf{0.014\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.011\scriptscriptstyle\pm\scriptstyle.001}$ $0.006\scriptscriptstyle\pm\scriptstyle.001$ $0.007\scriptscriptstyle\pm\scriptstyle.005$ $0.028\scriptscriptstyle\pm\scriptstyle.006$ $0.016\scriptscriptstyle\pm\scriptstyle.005$ $0.014\scriptscriptstyle\pm\scriptstyle.002$ $0.048\scriptscriptstyle\pm\scriptstyle.026$ $0.018\scriptscriptstyle\pm\scriptstyle.007$ $0.005\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.016\scriptscriptstyle\pm\scriptstyle.002}$
TVAE $0.089\scriptscriptstyle\pm\scriptstyle.004$ $0.193\scriptscriptstyle\pm\scriptstyle.001$ $0.018\scriptscriptstyle\pm\scriptstyle.001$ $0.012\scriptscriptstyle\pm\scriptstyle.006$ $0.050\scriptscriptstyle\pm\scriptstyle.010$ $0.055\scriptscriptstyle\pm\scriptstyle.010$ $0.040\scriptscriptstyle\pm\scriptstyle.005$ $0.058\scriptscriptstyle\pm\scriptstyle.000$ $0.055\scriptscriptstyle\pm\scriptstyle.006$ $0.084\scriptscriptstyle\pm\scriptstyle.001$ $0.072\scriptscriptstyle\pm\scriptstyle.009$
CTGAN $0.067\scriptscriptstyle\pm\scriptstyle.002$ $0.067\scriptscriptstyle\pm\scriptstyle.002$ $0.048\scriptscriptstyle\pm\scriptstyle.002$ $0.265\scriptscriptstyle\pm\scriptstyle.011$ $0.121\scriptscriptstyle\pm\scriptstyle.006$ $0.207\scriptscriptstyle\pm\scriptstyle.008$ $0.061\scriptscriptstyle\pm\scriptstyle.003$ $0.136\scriptscriptstyle\pm\scriptstyle.006$ $0.015\scriptscriptstyle\pm\scriptstyle.006$ $0.036\scriptscriptstyle\pm\scriptstyle.002$ $0.042\scriptscriptstyle\pm\scriptstyle.003$
TabDDPM $0.017\scriptscriptstyle\pm\scriptstyle.001$ $0.014\scriptscriptstyle\pm\scriptstyle.002$ $0.011\scriptscriptstyle\pm\scriptstyle.001$ $0.009\scriptscriptstyle\pm\scriptstyle.005$ $0.016\scriptscriptstyle\pm\scriptstyle.003$ $0.017\scriptscriptstyle\pm\scriptstyle.003$ $010$001 $0.009\scriptscriptstyle\pm\scriptstyle.003$ $0.013\scriptscriptstyle\pm\scriptstyle.006$ $0.006\scriptscriptstyle\pm\scriptstyle.001$ $0.018\scriptscriptstyle\pm\scriptstyle.006$
GReaT $0.056\scriptscriptstyle\pm\scriptstyle.003$ $0.044\scriptscriptstyle\pm\scriptstyle.002$ $0.058\scriptscriptstyle\pm\scriptstyle.001$ $0.005\scriptscriptstyle\pm\scriptstyle.004$ $0.034\scriptscriptstyle\pm\scriptstyle.008$ $0.016\scriptscriptstyle\pm\scriptstyle.001$ $0.038\scriptscriptstyle\pm\scriptstyle.004$ $0.054\scriptscriptstyle\pm\scriptstyle.004$ $0.018\scriptscriptstyle\pm\scriptstyle.007$ $0.108\scriptscriptstyle\pm\scriptstyle.006$
PGM ($\epsilon=1$) $\mathbf{0.020\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.019\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.009\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.039\scriptscriptstyle\pm\scriptstyle.006}$ $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.007}$ $\mathbf{0.026\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.030\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.016\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.032\scriptscriptstyle\pm\scriptstyle.004}$
PrivSyn ($\epsilon=1$) $0.028\scriptscriptstyle\pm\scriptstyle.002$ $0.041\scriptscriptstyle\pm\scriptstyle.006$ $020$002 $0.041\scriptscriptstyle\pm\scriptstyle.005$ $0.095\scriptscriptstyle\pm\scriptstyle.009$ $0.105\scriptscriptstyle\pm\scriptstyle.012$ $0.113\scriptscriptstyle\pm\scriptstyle.010$ $0.069\scriptscriptstyle\pm\scriptstyle.035$ $0.047\scriptscriptstyle\pm\scriptstyle.014$ $0.029\scriptscriptstyle\pm\scriptstyle.002$ $0.059\scriptscriptstyle\pm\scriptstyle.005$
PATE-GAN ($\epsilon=1$) $0.152\scriptscriptstyle\pm\scriptstyle.001$ $0.189\scriptscriptstyle\pm\scriptstyle.002$ $0.136\scriptscriptstyle\pm\scriptstyle.002$ $0.176\scriptscriptstyle\pm\scriptstyle.009$ $0.080\scriptscriptstyle\pm\scriptstyle.003$ $0.630\scriptscriptstyle\pm\scriptstyle.007$ $0.070\scriptscriptstyle\pm\scriptstyle.003$ $0.494\scriptscriptstyle\pm\scriptstyle.003$ $0.189\scriptscriptstyle\pm\scriptstyle.010$ $0.044\scriptscriptstyle\pm\scriptstyle.002$ $0.177\scriptscriptstyle\pm\scriptstyle.003$
TableDiffusion ($\epsilon=1$) $0.195\scriptscriptstyle\pm\scriptstyle.002$ $0.243\scriptscriptstyle\pm\scriptstyle.002$ $0.091\scriptscriptstyle\pm\scriptstyle.002$ $0.131\scriptscriptstyle\pm\scriptstyle.006$ $0.360\scriptscriptstyle\pm\scriptstyle.004$ $0.147\scriptscriptstyle\pm\scriptstyle.008$ $0.206\scriptscriptstyle\pm\scriptstyle.003$ $0.140\scriptscriptstyle\pm\scriptstyle.002$ $0.283\scriptscriptstyle\pm\scriptstyle.008$ $0.103\scriptscriptstyle\pm\scriptstyle.001$ $0.310\scriptscriptstyle\pm\scriptstyle.003$
Table D19: Fidelity evaluation of one-way numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.212\scriptscriptstyle\pm\scriptstyle.012$ $0.097\scriptscriptstyle\pm\scriptstyle.001$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $0.037\scriptscriptstyle\pm\scriptstyle.000$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $0.039\scriptscriptstyle\pm\scriptstyle.001$ $0.040\scriptscriptstyle\pm\scriptstyle.000$ $0.034\scriptscriptstyle\pm\scriptstyle.000$ $0.065\scriptscriptstyle\pm\scriptstyle.001$ $0.029\scriptscriptstyle\pm\scriptstyle.000$ $0.042\scriptscriptstyle\pm\scriptstyle.000$
PrivSyn $0.013\scriptscriptstyle\pm\scriptstyle.001$ $0.023\scriptscriptstyle\pm\scriptstyle.001$ $0.010\scriptscriptstyle\pm\scriptstyle.000$ $0.019\scriptscriptstyle\pm\scriptstyle.002$ $0.026\scriptscriptstyle\pm\scriptstyle.000$ $0.019\scriptscriptstyle\pm\scriptstyle.002$ $0.029\scriptscriptstyle\pm\scriptstyle.001$ $0.010\scriptscriptstyle\pm\scriptstyle.000$ $\mathbf{0.013\scriptscriptstyle\pm\scriptstyle.000}$ $0.023\scriptscriptstyle\pm\scriptstyle.001$ $0.009\scriptscriptstyle\pm\scriptstyle.000$
TVAE $0.024\scriptscriptstyle\pm\scriptstyle.000$ $0.014\scriptscriptstyle\pm\scriptstyle.001$ $0.014\scriptscriptstyle\pm\scriptstyle.000$ $0.020\scriptscriptstyle\pm\scriptstyle.001$ $0.016\scriptscriptstyle\pm\scriptstyle.001$ $0.026\scriptscriptstyle\pm\scriptstyle.002$ $0.021\scriptscriptstyle\pm\scriptstyle.001$ $0.017\scriptscriptstyle\pm\scriptstyle.001$ $0.018\scriptscriptstyle\pm\scriptstyle.000$ $0.024\scriptscriptstyle\pm\scriptstyle.004$ $0.025\scriptscriptstyle\pm\scriptstyle.000$
CTGAN $\mathbf{0.007\scriptscriptstyle\pm\scriptstyle.000}$ $\mathbf{0.008\scriptscriptstyle\pm\scriptstyle.000}$ $0.023\scriptscriptstyle\pm\scriptstyle.001$ $0.039\scriptscriptstyle\pm\scriptstyle.001$ $0.033\scriptscriptstyle\pm\scriptstyle.001$ $0.021\scriptscriptstyle\pm\scriptstyle.000$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $0.036\scriptscriptstyle\pm\scriptstyle.001$ $0.016\scriptscriptstyle\pm\scriptstyle.001$ $0.062\scriptscriptstyle\pm\scriptstyle.003$ $0.021\scriptscriptstyle\pm\scriptstyle.001$
TabDDPM $0.007\scriptscriptstyle\pm\scriptstyle.001$ $0.012\scriptscriptstyle\pm\scriptstyle.000$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.014\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.005\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.009\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.008\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.007\scriptscriptstyle\pm\scriptstyle.001}$ $0.037\scriptscriptstyle\pm\scriptstyle.000$ $0.013\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.007\scriptscriptstyle\pm\scriptstyle.000}$
GReaT $0.012\scriptscriptstyle\pm\scriptstyle.001$ $0.021\scriptscriptstyle\pm\scriptstyle.002$ $0.033\scriptscriptstyle\pm\scriptstyle.001$ $0.026\scriptscriptstyle\pm\scriptstyle.002$ $010$001 $0.032\scriptscriptstyle\pm\scriptstyle.002$ $0.026\scriptscriptstyle\pm\scriptstyle.001$ $0.011\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.010\scriptscriptstyle\pm\scriptstyle.002}$ $0.013\scriptscriptstyle\pm\scriptstyle.001$
PGM ($\epsilon=1$) $0.216\scriptscriptstyle\pm\scriptstyle.015$ $0.098\scriptscriptstyle\pm\scriptstyle.001$ $0.035\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.040\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.034\scriptscriptstyle\pm\scriptstyle.002}$ $0.043\scriptscriptstyle\pm\scriptstyle.000$ $\mathbf{0.041\scriptscriptstyle\pm\scriptstyle.000}$ $\mathbf{0.034\scriptscriptstyle\pm\scriptstyle.001}$ $0.067\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.031\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.043\scriptscriptstyle\pm\scriptstyle.000}$
PrivSyn ($\epsilon=1$) $\mathbf{0.020\scriptscriptstyle\pm\scriptstyle.001}$ $0.044\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.028\scriptscriptstyle\pm\scriptstyle.001}$ $0.168\scriptscriptstyle\pm\scriptstyle.004$ $0.072\scriptscriptstyle\pm\scriptstyle.003$ $0.098\scriptscriptstyle\pm\scriptstyle.003$ $0.129\scriptscriptstyle\pm\scriptstyle.004$ $0.093\scriptscriptstyle\pm\scriptstyle.004$ $0.056\scriptscriptstyle\pm\scriptstyle.001$ $0.078\scriptscriptstyle\pm\scriptstyle.002$ $0.103\scriptscriptstyle\pm\scriptstyle.003$
PATE-GAN ($\epsilon=1$) $0.028\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.037\scriptscriptstyle\pm\scriptstyle.001}$ $0.063\scriptscriptstyle\pm\scriptstyle.001$ $0.150\scriptscriptstyle\pm\scriptstyle.000$ $0.204\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.035\scriptscriptstyle\pm\scriptstyle.001}$ $0.230\scriptscriptstyle\pm\scriptstyle.000$ $0.236\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.040\scriptscriptstyle\pm\scriptstyle.000}$ $0.324\scriptscriptstyle\pm\scriptstyle.001$ $0.168\scriptscriptstyle\pm\scriptstyle.000$
TableDiffusion ($\epsilon=1$) $0.046\scriptscriptstyle\pm\scriptstyle.001$ $0.050\scriptscriptstyle\pm\scriptstyle.000$ $0.067\scriptscriptstyle\pm\scriptstyle.000$ $0.126\scriptscriptstyle\pm\scriptstyle.000$ $0.072\scriptscriptstyle\pm\scriptstyle.002$ $0.244\scriptscriptstyle\pm\scriptstyle.002$ $0.132\scriptscriptstyle\pm\scriptstyle.000$ $0.090\scriptscriptstyle\pm\scriptstyle.001$ $0.067\scriptscriptstyle\pm\scriptstyle.001$ $0.156\scriptscriptstyle\pm\scriptstyle.001$ $0.072\scriptscriptstyle\pm\scriptstyle.001$
Table D20: Fidelity evaluation of two-way categorical-categorical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.049\scriptscriptstyle\pm\scriptstyle.001$ $0.044\scriptscriptstyle\pm\scriptstyle.002$ $0.033\scriptscriptstyle\pm\scriptstyle.001$ $0.087\scriptscriptstyle\pm\scriptstyle.005$ $0.020\scriptscriptstyle\pm\scriptstyle.001$ $0.025\scriptscriptstyle\pm\scriptstyle.001$ $0.037\scriptscriptstyle\pm\scriptstyle.010$
PrivSyn $0.038\scriptscriptstyle\pm\scriptstyle.003$ $0.033\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.001}$ $0.151\scriptscriptstyle\pm\scriptstyle.016$ $0.038\scriptscriptstyle\pm\scriptstyle.003$ $0.022\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.035\scriptscriptstyle\pm\scriptstyle.005}$
TVAE $0.150\scriptscriptstyle\pm\scriptstyle.004$ $0.336\scriptscriptstyle\pm\scriptstyle.001$ $0.030\scriptscriptstyle\pm\scriptstyle.001$ $0.087\scriptscriptstyle\pm\scriptstyle.014$ $0.076\scriptscriptstyle\pm\scriptstyle.006$ $0.161\scriptscriptstyle\pm\scriptstyle.002$ $0.135\scriptscriptstyle\pm\scriptstyle.010$
CTGAN $0.127\scriptscriptstyle\pm\scriptstyle.001$ $0.138\scriptscriptstyle\pm\scriptstyle.001$ $0.077\scriptscriptstyle\pm\scriptstyle.003$ $0.259\scriptscriptstyle\pm\scriptstyle.006$ $0.126\scriptscriptstyle\pm\scriptstyle.003$ $0.062\scriptscriptstyle\pm\scriptstyle.002$ $0.1$005 -
TabDDPM $\mathbf{0.036\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.030\scriptscriptstyle\pm\scriptstyle.002}$ $0.019\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.032\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.019\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.010\scriptscriptstyle\pm\scriptstyle.001}$ $0.037\scriptscriptstyle\pm\scriptstyle.007$
GReaT $0.092\scriptscriptstyle\pm\scriptstyle.003$ $0.074\scriptscriptstyle\pm\scriptstyle.002$ $0.094\scriptscriptstyle\pm\scriptstyle.002$ $0.074\scriptscriptstyle\pm\scriptstyle.007$ $0.064\scriptscriptstyle\pm\scriptstyle.006$ $0.205\scriptscriptstyle\pm\scriptstyle.006$
PGM ($\epsilon=1$) $0.081\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.074\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.038\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.154\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.066\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.027\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.112\scriptscriptstyle\pm\scriptstyle.002}$
PrivSyn ($\epsilon=1$) $\mathbf{0.077\scriptscriptstyle\pm\scriptstyle.003}$ $0.118\scriptscriptstyle\pm\scriptstyle.007$ $0.046\scriptscriptstyle\pm\scriptstyle.002$ $0.267\scriptscriptstyle\pm\scriptstyle.005$ $0.205\scriptscriptstyle\pm\scriptstyle.016$ $0.066\scriptscriptstyle\pm\scriptstyle.003$ $0.143\scriptscriptstyle\pm\scriptstyle.011$
PATE-GAN ($\epsilon=1$) $0.278\scriptscriptstyle\pm\scriptstyle.002$ $0.347\scriptscriptstyle\pm\scriptstyle.001$ $0.210\scriptscriptstyle\pm\scriptstyle.002$ $0.253\scriptscriptstyle\pm\scriptstyle.006$ $0.129\scriptscriptstyle\pm\scriptstyle.005$ $0.073\scriptscriptstyle\pm\scriptstyle.003$ $0.273\scriptscriptstyle\pm\scriptstyle.005$
TableDiffusion ($\epsilon=1$) $0.323\scriptscriptstyle\pm\scriptstyle.002$ $0.361\scriptscriptstyle\pm\scriptstyle.003$ $0.155\scriptscriptstyle\pm\scriptstyle.002$ $0.620\scriptscriptstyle\pm\scriptstyle.002$ $0.348\scriptscriptstyle\pm\scriptstyle.004$ $0.196\scriptscriptstyle\pm\scriptstyle.001$ $0.466\scriptscriptstyle\pm\scriptstyle.003$
Table D21: Fidelity evaluation of two-way categorical-numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.230\scriptscriptstyle\pm\scriptstyle.012$ $0.110\scriptscriptstyle\pm\scriptstyle.002$ $0.038\scriptscriptstyle\pm\scriptstyle.002$ $0.058\scriptscriptstyle\pm\scriptstyle.002$ $0.045\scriptscriptstyle\pm\scriptstyle.003$ $0.051\scriptscriptstyle\pm\scriptstyle.001$ $0.049\scriptscriptstyle\pm\scriptstyle.004$ $040$004 $0.072\scriptscriptstyle\pm\scriptstyle.001$ $0.048\scriptscriptstyle\pm\scriptstyle.007$
PrivSyn $0.029\scriptscriptstyle\pm\scriptstyle.001$ $0.036\scriptscriptstyle\pm\scriptstyle.001$ $0.018\scriptscriptstyle\pm\scriptstyle.005$ $0.065\scriptscriptstyle\pm\scriptstyle.004$ $0.045\scriptscriptstyle\pm\scriptstyle.005$ $0.043\scriptscriptstyle\pm\scriptstyle.002$ $0$022 $0.032\scriptscriptstyle\pm\scriptstyle.007$ $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.001}$ $0.041\scriptscriptstyle\pm\scriptstyle.002$
TVAE $0.113\scriptscriptstyle\pm\scriptstyle.003$ $0.207\scriptscriptstyle\pm\scriptstyle.001$ $0.027\scriptscriptstyle\pm\scriptstyle.005$ $0.074\scriptscriptstyle\pm\scriptstyle.010$ $0.077\scriptscriptstyle\pm\scriptstyle.010$ $0.070\scriptscriptstyle\pm\scriptstyle.005$ $0.082\scriptscriptstyle\pm\scriptstyle.001$ $0.085\scriptscriptstyle\pm\scriptstyle.005$ $0.103\scriptscriptstyle\pm\scriptstyle.002$ $0.103\scriptscriptstyle\pm\scriptstyle.008$
CTGAN $0.077\scriptscriptstyle\pm\scriptstyle.002$ $0.077\scriptscriptstyle\pm\scriptstyle.002$ $0.288\scriptscriptstyle\pm\scriptstyle.011$ $0.165\scriptscriptstyle\pm\scriptstyle.007$ $0.246\scriptscriptstyle\pm\scriptstyle.009$ $0.092\scriptscriptstyle\pm\scriptstyle.003$ $0.175\scriptscriptstyle\pm\scriptstyle.006$ $0.059\scriptscriptstyle\pm\scriptstyle.006$ $0.053\scriptscriptstyle\pm\scriptstyle.001$ 0.111±.004plus-or-minus0.111.0040.$1$004 -
TabDDPM $\mathbf{0.025\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.027\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.006}$ $\mathbf{0.033\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.023\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.022\scriptscriptstyle\pm\scriptstyle.007}$ $0.043\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.035\scriptscriptstyle\pm\scriptstyle.006}$
GReaT $0.068\scriptscriptstyle\pm\scriptstyle.002$ $0.065\scriptscriptstyle\pm\scriptstyle.003$ $0.041\scriptscriptstyle\pm\scriptstyle.003$ $0.064\scriptscriptstyle\pm\scriptstyle.008$ $0.030\scriptscriptstyle\pm\scriptstyle.001$ $0.072\scriptscriptstyle\pm\scriptstyle.006$ $0.084\scriptscriptstyle\pm\scriptstyle.005$ $0.033\scriptscriptstyle\pm\scriptstyle.006$ $0.120\scriptscriptstyle\pm\scriptstyle.006$
PGM ($\epsilon=1$) $0.243\scriptscriptstyle\pm\scriptstyle.017$ $0.119\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.042\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.099\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.060\scriptscriptstyle\pm\scriptstyle.006}$ $\mathbf{0.081\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.080\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.059\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.075\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.088\scriptscriptstyle\pm\scriptstyle.001}$
PrivSyn ($\epsilon=1$) $\mathbf{0.055\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.089\scriptscriptstyle\pm\scriptstyle.005}$ $0.070\scriptscriptstyle\pm\scriptstyle.005$ $0.269\scriptscriptstyle\pm\scriptstyle.010$ $0.220\scriptscriptstyle\pm\scriptstyle.008$ $0.216\scriptscriptstyle\pm\scriptstyle.012$ $0.212\scriptscriptstyle\pm\scriptstyle.037$ $0.146\scriptscriptstyle\pm\scriptstyle.014$ $0.086\scriptscriptstyle\pm\scriptstyle.003$ $0.161\scriptscriptstyle\pm\scriptstyle.005$
PATE-GAN ($\epsilon=1$) $0.182\scriptscriptstyle\pm\scriptstyle.001$ $0.228\scriptscriptstyle\pm\scriptstyle.001$ $0.240\scriptscriptstyle\pm\scriptstyle.009$ $0.237\scriptscriptstyle\pm\scriptstyle.003$ $0.837\scriptscriptstyle\pm\scriptstyle.009$ $0.113\scriptscriptstyle\pm\scriptstyle.003$ $0.724\scriptscriptstyle\pm\scriptstyle.004$ $0.427\scriptscriptstyle\pm\scriptstyle.010$ $0.085\scriptscriptstyle\pm\scriptstyle.002$ $0.501\scriptscriptstyle\pm\scriptstyle.004$
TableDiffusion ($\epsilon=1$) $0.242\scriptscriptstyle\pm\scriptstyle.003$ $0.293\scriptscriptstyle\pm\scriptstyle.002$ $0.197\scriptscriptstyle\pm\scriptstyle.006$ $0.487\scriptscriptstyle\pm\scriptstyle.004$ $0.226\scriptscriptstyle\pm\scriptstyle.008$ $0.450\scriptscriptstyle\pm\scriptstyle.003$ $0.274\scriptscriptstyle\pm\scriptstyle.002$ $0.373\scriptscriptstyle\pm\scriptstyle.009$ $0.171\scriptscriptstyle\pm\scriptstyle.000$ $0.467\scriptscriptstyle\pm\scriptstyle.003$
Table D22: Fidelity evaluation of two-way numerical-numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.425\scriptscriptstyle\pm\scriptstyle.023$ $0.195\scriptscriptstyle\pm\scriptstyle.002$ $0.070\scriptscriptstyle\pm\scriptstyle.001$ $0.083\scriptscriptstyle\pm\scriptstyle.001$ $0.072\scriptscriptstyle\pm\scriptstyle.001$ $0.085\scriptscriptstyle\pm\scriptstyle.001$ $0.103\scriptscriptstyle\pm\scriptstyle.000$ $0.069\scriptscriptstyle\pm\scriptstyle.001$ $0.135\scriptscriptstyle\pm\scriptstyle.002$ $0.060\scriptscriptstyle\pm\scriptstyle.001$ $0.090\scriptscriptstyle\pm\scriptstyle.001$
PrivSyn $0.027\scriptscriptstyle\pm\scriptstyle.001$ $0.048\scriptscriptstyle\pm\scriptstyle.003$ $0.024\scriptscriptstyle\pm\scriptstyle.000$ $0.057\scriptscriptstyle\pm\scriptstyle.002$ $0.054\scriptscriptstyle\pm\scriptstyle.000$ $0.057\scriptscriptstyle\pm\scriptstyle.002$ $0.084\scriptscriptstyle\pm\scriptstyle.001$ $0.036\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.031\scriptscriptstyle\pm\scriptstyle.000}$ $0.052\scriptscriptstyle\pm\scriptstyle.001$ $0.026\scriptscriptstyle\pm\scriptstyle.000$
TVAE $0.047\scriptscriptstyle\pm\scriptstyle.001$ $0.028\scriptscriptstyle\pm\scriptstyle.001$ $0.032\scriptscriptstyle\pm\scriptstyle.001$ $0.045\scriptscriptstyle\pm\scriptstyle.002$ $0.039\scriptscriptstyle\pm\scriptstyle.002$ $0.061\scriptscriptstyle\pm\scriptstyle.002$ $0.050\scriptscriptstyle\pm\scriptstyle.001$ $0.037\scriptscriptstyle\pm\scriptstyle.001$ $0.037\scriptscriptstyle\pm\scriptstyle.001$ $0.055\scriptscriptstyle\pm\scriptstyle.007$ $0.052\scriptscriptstyle\pm\scriptstyle.000$
CTGAN $0.016\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.018\scriptscriptstyle\pm\scriptstyle.000}$ $0.051\scriptscriptstyle\pm\scriptstyle.002$ $0.084\scriptscriptstyle\pm\scriptstyle.001$ $0.072\scriptscriptstyle\pm\scriptstyle.001$ $0.058\scriptscriptstyle\pm\scriptstyle.001$ $0.080\scriptscriptstyle\pm\scriptstyle.001$ $0.088\scriptscriptstyle\pm\scriptstyle.002$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $0.134\scriptscriptstyle\pm\scriptstyle.005$ $0.044\scriptscriptstyle\pm\scriptstyle.001$
TabDDPM $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.001}$ $0.026\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.033\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.013\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.026\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.023\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.016\scriptscriptstyle\pm\scriptstyle.002}$ $0.076\scriptscriptstyle\pm\scriptstyle.001$ $0.035\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.001}$
GReaT $0.024\scriptscriptstyle\pm\scriptstyle.002$ $0.042\scriptscriptstyle\pm\scriptstyle.005$ $0.067\scriptscriptstyle\pm\scriptstyle.003$ $0.055\scriptscriptstyle\pm\scriptstyle.004$ $0.023\scriptscriptstyle\pm\scriptstyle.001$ $0.068\scriptscriptstyle\pm\scriptstyle.004$ $0.055\scriptscriptstyle\pm\scriptstyle.002$ $0.025\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.028\scriptscriptstyle\pm\scriptstyle.002}$ $0.026\scriptscriptstyle\pm\scriptstyle.002$
PGM ($\epsilon=1$) $0.432\scriptscriptstyle\pm\scriptstyle.030$ $0.203\scriptscriptstyle\pm\scriptstyle.003$ $0.084\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.098\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.079\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.099\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.109\scriptscriptstyle\pm\scriptstyle.000}$ $\mathbf{0.099\scriptscriptstyle\pm\scriptstyle.001}$ $0.137\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.090\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.092\scriptscriptstyle\pm\scriptstyle.001}$
PrivSyn ($\epsilon=1$) $\mathbf{0.043\scriptscriptstyle\pm\scriptstyle.002}$ $0.091\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.068\scriptscriptstyle\pm\scriptstyle.001}$ $0.340\scriptscriptstyle\pm\scriptstyle.009$ $0.171\scriptscriptstyle\pm\scriptstyle.005$ $0.205\scriptscriptstyle\pm\scriptstyle.005$ $0.268\scriptscriptstyle\pm\scriptstyle.007$ $0.221\scriptscriptstyle\pm\scriptstyle.006$ $0.115\scriptscriptstyle\pm\scriptstyle.002$ $0.176\scriptscriptstyle\pm\scriptstyle.005$ $0.209\scriptscriptstyle\pm\scriptstyle.005$
PATE-GAN ($\epsilon=1$) $0.056\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.078\scriptscriptstyle\pm\scriptstyle.002}$ $0.131\scriptscriptstyle\pm\scriptstyle.001$ $0.302\scriptscriptstyle\pm\scriptstyle.001$ $0.409\scriptscriptstyle\pm\scriptstyle.005$ $0.082\scriptscriptstyle\pm\scriptstyle.002$ $0.461\scriptscriptstyle\pm\scriptstyle.001$ $0.474\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.082\scriptscriptstyle\pm\scriptstyle.001}$ $0.649\scriptscriptstyle\pm\scriptstyle.002$ $0.335\scriptscriptstyle\pm\scriptstyle.000$
TableDiffusion ($\epsilon=1$) $0.092\scriptscriptstyle\pm\scriptstyle.001$ $0.099\scriptscriptstyle\pm\scriptstyle.001$ $0.133\scriptscriptstyle\pm\scriptstyle.001$ $0.252\scriptscriptstyle\pm\scriptstyle.001$ $0.147\scriptscriptstyle\pm\scriptstyle.003$ $0.488\scriptscriptstyle\pm\scriptstyle.004$ $0.265\scriptscriptstyle\pm\scriptstyle.001$ $0.181\scriptscriptstyle\pm\scriptstyle.002$ $0.136\scriptscriptstyle\pm\scriptstyle.001$ $0.314\scriptscriptstyle\pm\scriptstyle.001$ $0.145\scriptscriptstyle\pm\scriptstyle.001$
Table D23: Fidelity evaluation of one-way categorical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $\mathbf{0.013\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.023\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.012\scriptscriptstyle\pm\scriptstyle.001}$ $0.009\scriptscriptstyle\pm\scriptstyle.004$ $0.044\scriptscriptstyle\pm\scriptstyle.007$ $\mathbf{0.017\scriptscriptstyle\pm\scriptstyle.004}$ $0.031\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.012\scriptscriptstyle\pm\scriptstyle.007}$ $0.030\scriptscriptstyle\pm\scriptstyle.011$ $\mathbf{0.005\scriptscriptstyle\pm\scriptstyle.001}$ $0.062\scriptscriptstyle\pm\scriptstyle.012$
PrivSyn $0.015\scriptscriptstyle\pm\scriptstyle.001$ $0.023\scriptscriptstyle\pm\scriptstyle.004$ $0.012\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.007\scriptscriptstyle\pm\scriptstyle.004}$ $040$004 $0.018\scriptscriptstyle\pm\scriptstyle.006$ $0.031\scriptscriptstyle\pm\scriptstyle.007$ $0.016\scriptscriptstyle\pm\scriptstyle.005$ $\mathbf{0.020\scriptscriptstyle\pm\scriptstyle.005}$ $0.008\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.057\scriptscriptstyle\pm\scriptstyle.009}$
TVAE $0.090\scriptscriptstyle\pm\scriptstyle.003$ $0.191\scriptscriptstyle\pm\scriptstyle.003$ $0.021\scriptscriptstyle\pm\scriptstyle.001$ $0.009\scriptscriptstyle\pm\scriptstyle.005$ $0.083\scriptscriptstyle\pm\scriptstyle.024$ $0.060\scriptscriptstyle\pm\scriptstyle.007$ $0.036\scriptscriptstyle\pm\scriptstyle.003$ $0.062\scriptscriptstyle\pm\scriptstyle.012$ $0.049\scriptscriptstyle\pm\scriptstyle.009$ $0.084\scriptscriptstyle\pm\scriptstyle.001$ $0.115\scriptscriptstyle\pm\scriptstyle.009$
CTGAN $0.069\scriptscriptstyle\pm\scriptstyle.004$ $0.065\scriptscriptstyle\pm\scriptstyle.002$ $0.054\scriptscriptstyle\pm\scriptstyle.001$ $0.251\scriptscriptstyle\pm\scriptstyle.008$ $0.102\scriptscriptstyle\pm\scriptstyle.011$ $0.205\scriptscriptstyle\pm\scriptstyle.008$ $0.073\scriptscriptstyle\pm\scriptstyle.005$ $0.135\scriptscriptstyle\pm\scriptstyle.014$ $0.028\scriptscriptstyle\pm\scriptstyle.009$ $0.036\scriptscriptstyle\pm\scriptstyle.001$ $0.068\scriptscriptstyle\pm\scriptstyle.010$
TabDDPM $0.017\scriptscriptstyle\pm\scriptstyle.001$ $0.025\scriptscriptstyle\pm\scriptstyle.001$ $0.014\scriptscriptstyle\pm\scriptstyle.001$ $0.009\scriptscriptstyle\pm\scriptstyle.003$ $0.055\scriptscriptstyle\pm\scriptstyle.011$ $\mathbf{0.017\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.029\scriptscriptstyle\pm\scriptstyle.004}$ $0.022\scriptscriptstyle\pm\scriptstyle.006$ $0.033\scriptscriptstyle\pm\scriptstyle.022$ $0.006\scriptscriptstyle\pm\scriptstyle.001$ $0.059\scriptscriptstyle\pm\scriptstyle.010$
GReaT $0.057\scriptscriptstyle\pm\scriptstyle.002$ $0.053\scriptscriptstyle\pm\scriptstyle.004$ $0.055\scriptscriptstyle\pm\scriptstyle.001$ $0.008\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.031\scriptscriptstyle\pm\scriptstyle.007}$ $0.018\scriptscriptstyle\pm\scriptstyle.007$ $0.038\scriptscriptstyle\pm\scriptstyle.008$ $0.058\scriptscriptstyle\pm\scriptstyle.010$ $0.032\scriptscriptstyle\pm\scriptstyle.005$ $0.112\scriptscriptstyle\pm\scriptstyle.010$
PGM ($\epsilon=1$) $\mathbf{0.016\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.025\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.013\scriptscriptstyle\pm\scriptstyle.000}$ $0.011\scriptscriptstyle\pm\scriptstyle.007$ $\mathbf{0.066\scriptscriptstyle\pm\scriptstyle.007}$ $\mathbf{0.020\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.041\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.041\scriptscriptstyle\pm\scriptstyle.015}$ $0.051\scriptscriptstyle\pm\scriptstyle.016$ $0.007\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.053\scriptscriptstyle\pm\scriptstyle.006}$
PrivSyn ($\epsilon=1$) $0.032\scriptscriptstyle\pm\scriptstyle.003$ $0.051\scriptscriptstyle\pm\scriptstyle.002$ $0.018\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.003\scriptscriptstyle\pm\scriptstyle.002}$ $0.138\scriptscriptstyle\pm\scriptstyle.025$ $0.077\scriptscriptstyle\pm\scriptstyle.016$ $0.085\scriptscriptstyle\pm\scriptstyle.012$ $0.086\scriptscriptstyle\pm\scriptstyle.016$ $\mathbf{0.029\scriptscriptstyle\pm\scriptstyle.015}$ $\mathbf{0.012\scriptscriptstyle\pm\scriptstyle.001}$ $0.104\scriptscriptstyle\pm\scriptstyle.009$
PATE-GAN ($\epsilon=1$) $0.152\scriptscriptstyle\pm\scriptstyle.003$ $0.184\scriptscriptstyle\pm\scriptstyle.002$ $0.136\scriptscriptstyle\pm\scriptstyle.002$ $0.183\scriptscriptstyle\pm\scriptstyle.009$ $0.079\scriptscriptstyle\pm\scriptstyle.003$ $0.636\scriptscriptstyle\pm\scriptstyle.005$ $0.073\scriptscriptstyle\pm\scriptstyle.004$ $0.496\scriptscriptstyle\pm\scriptstyle.012$ $0.204\scriptscriptstyle\pm\scriptstyle.018$ $0.042\scriptscriptstyle\pm\scriptstyle.001$ $0.152\scriptscriptstyle\pm\scriptstyle.007$
TableDiffusion ($\epsilon=1$) $0.194\scriptscriptstyle\pm\scriptstyle.002$ $0.247\scriptscriptstyle\pm\scriptstyle.002$ $0.089\scriptscriptstyle\pm\scriptstyle.002$ $0.133\scriptscriptstyle\pm\scriptstyle.008$ $0.371\scriptscriptstyle\pm\scriptstyle.004$ $0.150\scriptscriptstyle\pm\scriptstyle.007$ $0.194\scriptscriptstyle\pm\scriptstyle.004$ $0.140\scriptscriptstyle\pm\scriptstyle.003$ $0.254\scriptscriptstyle\pm\scriptstyle.010$ $0.102\scriptscriptstyle\pm\scriptstyle.000$ $0.320\scriptscriptstyle\pm\scriptstyle.007$
Table D24: Fidelity evaluation of one-way numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.195\scriptscriptstyle\pm\scriptstyle.004$ $0.095\scriptscriptstyle\pm\scriptstyle.001$ $0.033\scriptscriptstyle\pm\scriptstyle.000$ $0.054\scriptscriptstyle\pm\scriptstyle.003$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $0.043\scriptscriptstyle\pm\scriptstyle.001$ $0.042\scriptscriptstyle\pm\scriptstyle.000$ $0.046\scriptscriptstyle\pm\scriptstyle.003$ $0.052\scriptscriptstyle\pm\scriptstyle.004$ $0.036\scriptscriptstyle\pm\scriptstyle.002$ $0.048\scriptscriptstyle\pm\scriptstyle.001$
PrivSyn $0.013\scriptscriptstyle\pm\scriptstyle.000$ $0.024\scriptscriptstyle\pm\scriptstyle.000$ $0.011\scriptscriptstyle\pm\scriptstyle.000$ $0.032\scriptscriptstyle\pm\scriptstyle.005$ $0.027\scriptscriptstyle\pm\scriptstyle.000$ $0.027\scriptscriptstyle\pm\scriptstyle.002$ $0.031\scriptscriptstyle\pm\scriptstyle.003$ $0.019\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.011\scriptscriptstyle\pm\scriptstyle.000}$ $0.031\scriptscriptstyle\pm\scriptstyle.003$ $0.012\scriptscriptstyle\pm\scriptstyle.001$
TVAE $0.025\scriptscriptstyle\pm\scriptstyle.000$ $0.013\scriptscriptstyle\pm\scriptstyle.001$ $0.014\scriptscriptstyle\pm\scriptstyle.000$ $0.028\scriptscriptstyle\pm\scriptstyle.003$ $0.017\scriptscriptstyle\pm\scriptstyle.001$ $0.033\scriptscriptstyle\pm\scriptstyle.002$ $0.024\scriptscriptstyle\pm\scriptstyle.001$ $0.016\scriptscriptstyle\pm\scriptstyle.001$ $0.017\scriptscriptstyle\pm\scriptstyle.001$ $0.033\scriptscriptstyle\pm\scriptstyle.005$ $0.029\scriptscriptstyle\pm\scriptstyle.001$
CTGAN $0.009\scriptscriptstyle\pm\scriptstyle.000$ $\mathbf{0.009\scriptscriptstyle\pm\scriptstyle.001}$ $0.024\scriptscriptstyle\pm\scriptstyle.000$ $0.047\scriptscriptstyle\pm\scriptstyle.002$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $0.027\scriptscriptstyle\pm\scriptstyle.001$ $0.037\scriptscriptstyle\pm\scriptstyle.003$ $0.046\scriptscriptstyle\pm\scriptstyle.001$ $0.016\scriptscriptstyle\pm\scriptstyle.000$ $0.072\scriptscriptstyle\pm\scriptstyle.003$ $0.022\scriptscriptstyle\pm\scriptstyle.001$
TabDDPM $\mathbf{0.008\scriptscriptstyle\pm\scriptstyle.000}$ $0.012\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.013\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.017\scriptscriptstyle\pm\scriptstyle.004}$ $0.034\scriptscriptstyle\pm\scriptstyle.000$ $0.024\scriptscriptstyle\pm\scriptstyle.006$ $\mathbf{0.010\scriptscriptstyle\pm\scriptstyle.001}$
GReaT $0.013\scriptscriptstyle\pm\scriptstyle.000$ $0.021\scriptscriptstyle\pm\scriptstyle.002$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $0.036\scriptscriptstyle\pm\scriptstyle.002$ $0.009\scriptscriptstyle\pm\scriptstyle.000$ $0.039\scriptscriptstyle\pm\scriptstyle.004$ $0.027\scriptscriptstyle\pm\scriptstyle.002$ $0.019\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.020\scriptscriptstyle\pm\scriptstyle.006}$ $0.015\scriptscriptstyle\pm\scriptstyle.001$
PGM ($\epsilon=1$) $0.196\scriptscriptstyle\pm\scriptstyle.004$ $0.094\scriptscriptstyle\pm\scriptstyle.000$ $0.033\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.055\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.034\scriptscriptstyle\pm\scriptstyle.001}$ $0.045\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.043\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.046\scriptscriptstyle\pm\scriptstyle.002}$ $0.053\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.037\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.047\scriptscriptstyle\pm\scriptstyle.001}$
PrivSyn ($\epsilon=1$) $\mathbf{0.024\scriptscriptstyle\pm\scriptstyle.000}$ $0.044\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.026\scriptscriptstyle\pm\scriptstyle.002}$ $0.146\scriptscriptstyle\pm\scriptstyle.006$ $0.069\scriptscriptstyle\pm\scriptstyle.004$ $0.106\scriptscriptstyle\pm\scriptstyle.005$ $0.128\scriptscriptstyle\pm\scriptstyle.004$ $0.086\scriptscriptstyle\pm\scriptstyle.002$ $0.040\scriptscriptstyle\pm\scriptstyle.001$ $0.079\scriptscriptstyle\pm\scriptstyle.007$ $0.085\scriptscriptstyle\pm\scriptstyle.003$
PATE-GAN ($\epsilon=1$) $0.027\scriptscriptstyle\pm\scriptstyle.000$ $\mathbf{0.036\scriptscriptstyle\pm\scriptstyle.001}$ $0.063\scriptscriptstyle\pm\scriptstyle.001$ $0.166\scriptscriptstyle\pm\scriptstyle.003$ $0.204\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.033\scriptscriptstyle\pm\scriptstyle.003}$ $0.231\scriptscriptstyle\pm\scriptstyle.001$ $0.250\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.037\scriptscriptstyle\pm\scriptstyle.002}$ $0.341\scriptscriptstyle\pm\scriptstyle.009$ $0.167\scriptscriptstyle\pm\scriptstyle.002$
TableDiffusion ($\epsilon=1$) $0.046\scriptscriptstyle\pm\scriptstyle.000$ $0.048\scriptscriptstyle\pm\scriptstyle.000$ $0.066\scriptscriptstyle\pm\scriptstyle.001$ $0.146\scriptscriptstyle\pm\scriptstyle.001$ $0.072\scriptscriptstyle\pm\scriptstyle.001$ $0.243\scriptscriptstyle\pm\scriptstyle.002$ $0.132\scriptscriptstyle\pm\scriptstyle.000$ $0.099\scriptscriptstyle\pm\scriptstyle.002$ $0.060\scriptscriptstyle\pm\scriptstyle.001$ $0.174\scriptscriptstyle\pm\scriptstyle.005$ $0.065\scriptscriptstyle\pm\scriptstyle.001$
Table D25: Fidelity evaluation of two-way categorical-categorical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.048\scriptscriptstyle\pm\scriptstyle.003$ $0.063\scriptscriptstyle\pm\scriptstyle.002$ $0.041\scriptscriptstyle\pm\scriptstyle.001$ $0.122\scriptscriptstyle\pm\scriptstyle.005$ $0.059\scriptscriptstyle\pm\scriptstyle.004$ $0.025\scriptscriptstyle\pm\scriptstyle.001$ $0.122\scriptscriptstyle\pm\scriptstyle.021$
PrivSyn $\mathbf{0.038\scriptscriptstyle\pm\scriptstyle.001}$ $0.058\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.023\scriptscriptstyle\pm\scriptstyle.002}$ $0.150\scriptscriptstyle\pm\scriptstyle.010$ $0.064\scriptscriptstyle\pm\scriptstyle.009$ $0.024\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.107\scriptscriptstyle\pm\scriptstyle.008}$
TVAE $0.153\scriptscriptstyle\pm\scriptstyle.003$ 0.333±.003plus-or-minus0.333.0030.$3$003 $0.035\scriptscriptstyle\pm\scriptstyle.002$ $0.118\scriptscriptstyle\pm\scriptstyle.025$ $0.072\scriptscriptstyle\pm\scriptstyle.002$ $0.161\scriptscriptstyle\pm\scriptstyle.001$ $0.203\scriptscriptstyle\pm\scriptstyle.012$
CTGAN $0.128\scriptscriptstyle\pm\scriptstyle.005$ $0.137\scriptscriptstyle\pm\scriptstyle.002$ $0.083\scriptscriptstyle\pm\scriptstyle.001$ $0.256\scriptscriptstyle\pm\scriptstyle.005$ $0.147\scriptscriptstyle\pm\scriptstyle.008$ $0.063\scriptscriptstyle\pm\scriptstyle.002$ $0.132\scriptscriptstyle\pm\scriptstyle.009$
TabDDPM $0.039\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.054\scriptscriptstyle\pm\scriptstyle.001}$ $0.025\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.083\scriptscriptstyle\pm\scriptstyle.010}$ $\mathbf{0.055\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.010\scriptscriptstyle\pm\scriptstyle.002}$ $0.113\scriptscriptstyle\pm\scriptstyle.010$
GReaT $0.094\scriptscriptstyle\pm\scriptstyle.002$ $0.091\scriptscriptstyle\pm\scriptstyle.005$ $0.089\scriptscriptstyle\pm\scriptstyle.002$ $0.092\scriptscriptstyle\pm\scriptstyle.017$ $0.067\scriptscriptstyle\pm\scriptstyle.011$ $0.203\scriptscriptstyle\pm\scriptstyle.012$
PGM ($\epsilon=1$) $\mathbf{0.074\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.083\scriptscriptstyle\pm\scriptstyle.002}$ $0.043\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.172\scriptscriptstyle\pm\scriptstyle.007}$ $\mathbf{0.085\scriptscriptstyle\pm\scriptstyle.006}$ $\mathbf{0.027\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.144\scriptscriptstyle\pm\scriptstyle.010}$
PrivSyn ($\epsilon=1$) $0.082\scriptscriptstyle\pm\scriptstyle.004$ $0.135\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.041\scriptscriptstyle\pm\scriptstyle.003}$ $0.310\scriptscriptstyle\pm\scriptstyle.019$ $0.166\scriptscriptstyle\pm\scriptstyle.022$ $0.031\scriptscriptstyle\pm\scriptstyle.002$ $0.190\scriptscriptstyle\pm\scriptstyle.018$
PATE-GAN ($\epsilon=1$) $0.278\scriptscriptstyle\pm\scriptstyle.003$ $0.342\scriptscriptstyle\pm\scriptstyle.004$ $0.211\scriptscriptstyle\pm\scriptstyle.002$ $0.247\scriptscriptstyle\pm\scriptstyle.010$ $0.140\scriptscriptstyle\pm\scriptstyle.004$ $0.070\scriptscriptstyle\pm\scriptstyle.001$ $0.258\scriptscriptstyle\pm\scriptstyle.012$
TableDiffusion($\epsilon=1$) $0.321\scriptscriptstyle\pm\scriptstyle.002$ $0.365\scriptscriptstyle\pm\scriptstyle.003$ $0.153\scriptscriptstyle\pm\scriptstyle.002$ $0.620\scriptscriptstyle\pm\scriptstyle.007$ $0.332\scriptscriptstyle\pm\scriptstyle.006$ $0.195\scriptscriptstyle\pm\scriptstyle.000$ $0.494\scriptscriptstyle\pm\scriptstyle.008$
Table D26: Fidelity evaluation of two-way categorical-numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.212\scriptscriptstyle\pm\scriptstyle.005$ $0.119\scriptscriptstyle\pm\scriptstyle.002$ $0.043\scriptscriptstyle\pm\scriptstyle.004$ $0.106\scriptscriptstyle\pm\scriptstyle.007$ $0.053\scriptscriptstyle\pm\scriptstyle.005$ $0.080\scriptscriptstyle\pm\scriptstyle.004$ $0.058\scriptscriptstyle\pm\scriptstyle.006$ $0.077\scriptscriptstyle\pm\scriptstyle.012$ $0.060\scriptscriptstyle\pm\scriptstyle.005$ $0.106\scriptscriptstyle\pm\scriptstyle.010$
PrivSyn $0.029\scriptscriptstyle\pm\scriptstyle.001$ $0.049\scriptscriptstyle\pm\scriptstyle.003$ $0.019\scriptscriptstyle\pm\scriptstyle.004$ $0.090\scriptscriptstyle\pm\scriptstyle.005$ $0.048\scriptscriptstyle\pm\scriptstyle.006$ $0.069\scriptscriptstyle\pm\scriptstyle.006$ $0.079\scriptscriptstyle\pm\scriptstyle.006$ $0.042\scriptscriptstyle\pm\scriptstyle.006$ $\mathbf{0.020\scriptscriptstyle\pm\scriptstyle.001}$ $0.097\scriptscriptstyle\pm\scriptstyle.008$
TVAE $0.115\scriptscriptstyle\pm\scriptstyle.003$ $0.205\scriptscriptstyle\pm\scriptstyle.003$ $0.025\scriptscriptstyle\pm\scriptstyle.005$ $0.115\scriptscriptstyle\pm\scriptstyle.023$ $0.082\scriptscriptstyle\pm\scriptstyle.007$ $0.075\scriptscriptstyle\pm\scriptstyle.004$ $0.092\scriptscriptstyle\pm\scriptstyle.011$ $0.081\scriptscriptstyle\pm\scriptstyle.007$ $0.1$001 $0.158\scriptscriptstyle\pm\scriptstyle.011$
CTGAN $0.080\scriptscriptstyle\pm\scriptstyle.003$ $0.076\scriptscriptstyle\pm\scriptstyle.002$ $0.275\scriptscriptstyle\pm\scriptstyle.008$ $0.157\scriptscriptstyle\pm\scriptstyle.012$ $0.244\scriptscriptstyle\pm\scriptstyle.009$ $0.109\scriptscriptstyle\pm\scriptstyle.005$ $0.178\scriptscriptstyle\pm\scriptstyle.015$ $0.081\scriptscriptstyle\pm\scriptstyle.008$ $0.053\scriptscriptstyle\pm\scriptstyle.001$ $0.150\scriptscriptstyle\pm\scriptstyle.009$
TabDDPM $\mathbf{0.027\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.039\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.016\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.080\scriptscriptstyle\pm\scriptstyle.012}$ $\mathbf{0.025\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.055\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.041\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.051\scriptscriptstyle\pm\scriptstyle.024}$ $0.040\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.090\scriptscriptstyle\pm\scriptstyle.009}$
GReaT $0.070\scriptscriptstyle\pm\scriptstyle.002$ $0.074\scriptscriptstyle\pm\scriptstyle.004$ $0.045\scriptscriptstyle\pm\scriptstyle.004$ $0.078\scriptscriptstyle\pm\scriptstyle.006$ $0.032\scriptscriptstyle\pm\scriptstyle.006$ $0.081\scriptscriptstyle\pm\scriptstyle.008$ $0.089\scriptscriptstyle\pm\scriptstyle.010$ $0.055\scriptscriptstyle\pm\scriptstyle.004$ $0.138\scriptscriptstyle\pm\scriptstyle.012$
PGM ($\epsilon=1$) $0.219\scriptscriptstyle\pm\scriptstyle.005$ $0.121\scriptscriptstyle\pm\scriptstyle.002$ $0.044\scriptscriptstyle\pm\scriptstyle.007$ $\mathbf{0.139\scriptscriptstyle\pm\scriptstyle.006}$ $\mathbf{0.059\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.098\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.093\scriptscriptstyle\pm\scriptstyle.013}$ $\mathbf{0.102\scriptscriptstyle\pm\scriptstyle.016}$ $0.061\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.117\scriptscriptstyle\pm\scriptstyle.004}$
PrivSyn ($\epsilon=1$) $\mathbf{0.060\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.098\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.033\scriptscriptstyle\pm\scriptstyle.002}$ $0.290\scriptscriptstyle\pm\scriptstyle.022$ $0.201\scriptscriptstyle\pm\scriptstyle.015$ $0.197\scriptscriptstyle\pm\scriptstyle.017$ $0.226\scriptscriptstyle\pm\scriptstyle.017$ $0.124\scriptscriptstyle\pm\scriptstyle.017$ $\mathbf{0.053\scriptscriptstyle\pm\scriptstyle.001}$ $0.204\scriptscriptstyle\pm\scriptstyle.010$
PATE-GAN ($\epsilon=1$) $0.182\scriptscriptstyle\pm\scriptstyle.002$ $0.221\scriptscriptstyle\pm\scriptstyle.002$ $0.246\scriptscriptstyle\pm\scriptstyle.009$ $0.251\scriptscriptstyle\pm\scriptstyle.002$ $0.842\scriptscriptstyle\pm\scriptstyle.006$ $0.115\scriptscriptstyle\pm\scriptstyle.004$ $0.726\scriptscriptstyle\pm\scriptstyle.012$ $0.455\scriptscriptstyle\pm\scriptstyle.018$ $0.080\scriptscriptstyle\pm\scriptstyle.003$ $0.494\scriptscriptstyle\pm\scriptstyle.011$
TableDiffusion ($\epsilon=1$) $0.242\scriptscriptstyle\pm\scriptstyle.003$ $0.295\scriptscriptstyle\pm\scriptstyle.002$ $0.199\scriptscriptstyle\pm\scriptstyle.008$ $0.519\scriptscriptstyle\pm\scriptstyle.005$ $0.229\scriptscriptstyle\pm\scriptstyle.007$ $0.437\scriptscriptstyle\pm\scriptstyle.006$ $0.275\scriptscriptstyle\pm\scriptstyle.004$ $0.353\scriptscriptstyle\pm\scriptstyle.012$ $0.163\scriptscriptstyle\pm\scriptstyle.000$ $0.495\scriptscriptstyle\pm\scriptstyle.008$
Table D27: Fidelity evaluation of two-way numerical-numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.392\scriptscriptstyle\pm\scriptstyle.008$ $0.191\scriptscriptstyle\pm\scriptstyle.002$ $0.069\scriptscriptstyle\pm\scriptstyle.000$ $0.120\scriptscriptstyle\pm\scriptstyle.006$ $0.073\scriptscriptstyle\pm\scriptstyle.001$ $0.097\scriptscriptstyle\pm\scriptstyle.002$ $0.107\scriptscriptstyle\pm\scriptstyle.001$ $0.094\scriptscriptstyle\pm\scriptstyle.006$ $0.108\scriptscriptstyle\pm\scriptstyle.009$ $0.088\scriptscriptstyle\pm\scriptstyle.003$ $0.1$002
PrivSyn $0.027\scriptscriptstyle\pm\scriptstyle.000$ $0.050\scriptscriptstyle\pm\scriptstyle.001$ $0.026\scriptscriptstyle\pm\scriptstyle.001$ $0.083\scriptscriptstyle\pm\scriptstyle.009$ $0.056\scriptscriptstyle\pm\scriptstyle.001$ $0.073\scriptscriptstyle\pm\scriptstyle.002$ $0.089\scriptscriptstyle\pm\scriptstyle.003$ $0.049\scriptscriptstyle\pm\scriptstyle.006$ $\mathbf{0.026\scriptscriptstyle\pm\scriptstyle.001}$ $0.079\scriptscriptstyle\pm\scriptstyle.004$ $0.032\scriptscriptstyle\pm\scriptstyle.001$
TVAE $0.049\scriptscriptstyle\pm\scriptstyle.001$ $0.026\scriptscriptstyle\pm\scriptstyle.001$ $0.033\scriptscriptstyle\pm\scriptstyle.001$ $0.064\scriptscriptstyle\pm\scriptstyle.005$ $0.041\scriptscriptstyle\pm\scriptstyle.001$ $0.078\scriptscriptstyle\pm\scriptstyle.003$ $0.058\scriptscriptstyle\pm\scriptstyle.002$ $0.037\scriptscriptstyle\pm\scriptstyle.002$ $0.034\scriptscriptstyle\pm\scriptstyle.001$ $0.082\scriptscriptstyle\pm\scriptstyle.007$ $0.061\scriptscriptstyle\pm\scriptstyle.001$
CTGAN $0.018\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.019\scriptscriptstyle\pm\scriptstyle.002}$ $0.052\scriptscriptstyle\pm\scriptstyle.001$ $0.103\scriptscriptstyle\pm\scriptstyle.004$ $0.074\scriptscriptstyle\pm\scriptstyle.002$ $0.069\scriptscriptstyle\pm\scriptstyle.001$ $0.087\scriptscriptstyle\pm\scriptstyle.004$ $0.109\scriptscriptstyle\pm\scriptstyle.002$ $0.033\scriptscriptstyle\pm\scriptstyle.001$ $0.157\scriptscriptstyle\pm\scriptstyle.008$ $0.047\scriptscriptstyle\pm\scriptstyle.002$
TabDDPM $\mathbf{0.017\scriptscriptstyle\pm\scriptstyle.000}$ $0.025\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.049\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.015\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.054\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.035\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.035\scriptscriptstyle\pm\scriptstyle.006}$ $0.070\scriptscriptstyle\pm\scriptstyle.001$ $0.064\scriptscriptstyle\pm\scriptstyle.008$ $\mathbf{0.023\scriptscriptstyle\pm\scriptstyle.001}$
GReaT $0.026\scriptscriptstyle\pm\scriptstyle.001$ $0.042\scriptscriptstyle\pm\scriptstyle.003$ $0.068\scriptscriptstyle\pm\scriptstyle.002$ $0.079\scriptscriptstyle\pm\scriptstyle.003$ $0.023\scriptscriptstyle\pm\scriptstyle.001$ $0.086\scriptscriptstyle\pm\scriptstyle.008$ $0.059\scriptscriptstyle\pm\scriptstyle.004$ $0.040\scriptscriptstyle\pm\scriptstyle.005$ $\mathbf{0.060\scriptscriptstyle\pm\scriptstyle.008}$ $0.033\scriptscriptstyle\pm\scriptstyle.002$
PGM ($\epsilon=1$) $0.393\scriptscriptstyle\pm\scriptstyle.009$ $0.194\scriptscriptstyle\pm\scriptstyle.000$ $0.080\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.130\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.079\scriptscriptstyle\pm\scriptstyle.001}$ $0.108\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.113\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.121\scriptscriptstyle\pm\scriptstyle.004}$ $0.109\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.114\scriptscriptstyle\pm\scriptstyle.005}$ $\mathbf{0.102\scriptscriptstyle\pm\scriptstyle.001}$
PrivSyn ($\epsilon=1$) $\mathbf{0.049\scriptscriptstyle\pm\scriptstyle.001}$ $0.091\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.062\scriptscriptstyle\pm\scriptstyle.003}$ $0.297\scriptscriptstyle\pm\scriptstyle.012$ $0.167\scriptscriptstyle\pm\scriptstyle.007$ $0.219\scriptscriptstyle\pm\scriptstyle.010$ $0.268\scriptscriptstyle\pm\scriptstyle.008$ $0.205\scriptscriptstyle\pm\scriptstyle.007$ $0.083\scriptscriptstyle\pm\scriptstyle.002$ $0.181\scriptscriptstyle\pm\scriptstyle.010$ $0.174\scriptscriptstyle\pm\scriptstyle.006$
PATE-GAN ($\epsilon=1$) $0.055\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.074\scriptscriptstyle\pm\scriptstyle.001}$ $0.130\scriptscriptstyle\pm\scriptstyle.001$ $0.334\scriptscriptstyle\pm\scriptstyle.005$ $0.408\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.083\scriptscriptstyle\pm\scriptstyle.005}$ $0.461\scriptscriptstyle\pm\scriptstyle.002$ $0.501\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.077\scriptscriptstyle\pm\scriptstyle.004}$ $0.682\scriptscriptstyle\pm\scriptstyle.017$ $0.334\scriptscriptstyle\pm\scriptstyle.003$
TableDiffusion ($\epsilon=1$) $0.093\scriptscriptstyle\pm\scriptstyle.001$ $0.096\scriptscriptstyle\pm\scriptstyle.001$ $0.132\scriptscriptstyle\pm\scriptstyle.002$ $0.294\scriptscriptstyle\pm\scriptstyle.002$ $0.147\scriptscriptstyle\pm\scriptstyle.002$ $0.486\scriptscriptstyle\pm\scriptstyle.005$ $0.265\scriptscriptstyle\pm\scriptstyle.001$ $0$005 $0.121\scriptscriptstyle\pm\scriptstyle.001$ $0.348\scriptscriptstyle\pm\scriptstyle.011$ $0.132\scriptscriptstyle\pm\scriptstyle.001$
D.2 Utility Results
Here we show the utility results for all synthesizers. Specifically, the machine learning affinity results are shown in Table D28, and the query results are detailed in Table D29.
Table D28: Machine learning affinity of data synthesis. The privacy budget $\epsilon$ of HP synthesizers is set as $\infty$ (the top part), and the budget for DP synthesizers is set as $1$ (the bottom part). The lower value means the performance of synthetic data is more similar to real data. The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.086\scriptscriptstyle\pm\scriptstyle.001$ $0.193\scriptscriptstyle\pm\scriptstyle.002$ $0.037\scriptscriptstyle\pm\scriptstyle.003$ $0.073\scriptscriptstyle\pm\scriptstyle.001$ $0.255\scriptscriptstyle\pm\scriptstyle.002$ $0.035\scriptscriptstyle\pm\scriptstyle.003$ $0.332\scriptscriptstyle\pm\scriptstyle.001$ $0.146\scriptscriptstyle\pm\scriptstyle.002$ $0.096\scriptscriptstyle\pm\scriptstyle.003$ $0.498\scriptscriptstyle\pm\scriptstyle.001$ $0.270\scriptscriptstyle\pm\scriptstyle.002$ $0.347\scriptscriptstyle\pm\scriptstyle.003$
PrivSyn $0.120\scriptscriptstyle\pm\scriptstyle.003$ $0.040\scriptscriptstyle\pm\scriptstyle.001$ $0.057\scriptscriptstyle\pm\scriptstyle.002$ $0.085\scriptscriptstyle\pm\scriptstyle.003$ $0.532\scriptscriptstyle\pm\scriptstyle.001$ $0.039\scriptscriptstyle\pm\scriptstyle.002$ $0.604\scriptscriptstyle\pm\scriptstyle.003$ $0.406\scriptscriptstyle\pm\scriptstyle.001$ $0.210\scriptscriptstyle\pm\scriptstyle.002$ $1.992\scriptscriptstyle\pm\scriptstyle.003$ $0.518\scriptscriptstyle\pm\scriptstyle.001$ $0.201\scriptscriptstyle\pm\scriptstyle.002$
TVAE $0.035\scriptscriptstyle\pm\scriptstyle.002$ $0.011\scriptscriptstyle\pm\scriptstyle.003$ $0.031\scriptscriptstyle\pm\scriptstyle.001$ $0.075\scriptscriptstyle\pm\scriptstyle.002$ $0.217\scriptscriptstyle\pm\scriptstyle.003$ $0.059\scriptscriptstyle\pm\scriptstyle.001$ $0.294\scriptscriptstyle\pm\scriptstyle.002$ $0.128\scriptscriptstyle\pm\scriptstyle.003$ $0.245\scriptscriptstyle\pm\scriptstyle.001$ $0.147\scriptscriptstyle\pm\scriptstyle.002$ $0.336\scriptscriptstyle\pm\scriptstyle.003$ $0.091\scriptscriptstyle\pm\scriptstyle.001$
[…]
TabDDPM $0.014\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.007\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.007\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.085\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.003\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.011\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.043\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.047\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.140\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.047\scriptscriptstyle\pm\scriptstyle.003}$
GReaT $\mathbf{0.009\scriptscriptstyle\pm\scriptstyle.002}$ $0.009\scriptscriptstyle\pm\scriptstyle.003$ $0.020\scriptscriptstyle\pm\scriptstyle.001$ $0.033\scriptscriptstyle\pm\scriptstyle.002$ $0.183\scriptscriptstyle\pm\scriptstyle.003$ $0.017\scriptscriptstyle\pm\scriptstyle.001$ $0.117\scriptscriptstyle\pm\scriptstyle.002$ $0.050\scriptscriptstyle\pm\scriptstyle.003$ $0.038\scriptscriptstyle\pm\scriptstyle.001$ $0.292\scriptscriptstyle\pm\scriptstyle.002$ $0.083\scriptscriptstyle\pm\scriptstyle.003$
PGM ($\epsilon=1$) $\mathbf{0.101\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.048\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.041\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.093\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.489\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.054\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.531\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.245\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.241\scriptscriptstyle\pm\scriptstyle.002}$ $1.072\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{1.366\scriptscriptstyle\pm\scriptstyle.001}$ $0.340\scriptscriptstyle\pm\scriptstyle.002$
PrivSyn ($\epsilon=1$) $0.120\scriptscriptstyle\pm\scriptstyle.002$ $0.177\scriptscriptstyle\pm\scriptstyle.003$ $0.085\scriptscriptstyle\pm\scriptstyle.001$ $0.217\scriptscriptstyle\pm\scriptstyle.002$ $0.753\scriptscriptstyle\pm\scriptstyle.003$ $0.466\scriptscriptstyle\pm\scriptstyle.001$ $0.821\scriptscriptstyle\pm\scriptstyle.002$ $0.608\scriptscriptstyle\pm\scriptstyle.003$ $0.624\scriptscriptstyle\pm\scriptstyle.001$ $4.538\scriptscriptstyle\pm\scriptstyle.002$ $1.878\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.302\scriptscriptstyle\pm\scriptstyle.001}$
PATE-GAN ($\epsilon=1$) $0.126\scriptscriptstyle\pm\scriptstyle.001$ $0.135\scriptscriptstyle\pm\scriptstyle.002$ $0.530\scriptscriptstyle\pm\scriptstyle.003$ $0.394\scriptscriptstyle\pm\scriptstyle.001$ $0.781\scriptscriptstyle\pm\scriptstyle.002$ $0.781\scriptscriptstyle\pm\scriptstyle.003$ $0.877\scriptscriptstyle\pm\scriptstyle.001$ $0.755\scriptscriptstyle\pm\scriptstyle.002$ $2.119\scriptscriptstyle\pm\scriptstyle.003$ $0.259\scriptscriptstyle\pm\scriptstyle.001$ $2.325\scriptscriptstyle\pm\scriptstyle.002$ $0.405\scriptscriptstyle\pm\scriptstyle.003$
TableDiffusion ($\epsilon=1$) $0.198\scriptscriptstyle\pm\scriptstyle.002$ $0.135\scriptscriptstyle\pm\scriptstyle.003$ $0.074\scriptscriptstyle\pm\scriptstyle.001$ $0.133\scriptscriptstyle\pm\scriptstyle.002$ $0.904\scriptscriptstyle\pm\scriptstyle.003$ $0.981\scriptscriptstyle\pm\scriptstyle.001$ $0.968\scriptscriptstyle\pm\scriptstyle.002$ $0.439\scriptscriptstyle\pm\scriptstyle.003$ $0.287\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.781\scriptscriptstyle\pm\scriptstyle.002}$ $2.503\scriptscriptstyle\pm\scriptstyle.003$ $0.489\scriptscriptstyle\pm\scriptstyle.001$
Table D29: Query error of data synthesis. The privacy budget $\epsilon$ of HP synthesizers is set as $\infty$ (the top part), and the budget for DP synthesizers is set as $1$ (the bottom part). The best result of each category is in bold.
Adult Shoppers Phishing Magic Faults Bean Obesity Robot Abalone News Insurance Wine
PGM $0.056\scriptscriptstyle\pm\scriptstyle.018$ $0.044\scriptscriptstyle\pm\scriptstyle.005$ $\mathbf{0.009\scriptscriptstyle\pm\scriptstyle.001}$ $0.035\scriptscriptstyle\pm\scriptstyle.004$ $0.041\scriptscriptstyle\pm\scriptstyle.003$ $0.036\scriptscriptstyle\pm\scriptstyle.003$ $0.035\scriptscriptstyle\pm\scriptstyle.007$ $0.049\scriptscriptstyle\pm\scriptstyle.005$ $040$004 $0.033\scriptscriptstyle\pm\scriptstyle.005$ $0.039\scriptscriptstyle\pm\scriptstyle.004$ $0.042\scriptscriptstyle\pm\scriptstyle.005$
PrivSyn $0.009\scriptscriptstyle\pm\scriptstyle.002$ $0.011\scriptscriptstyle\pm\scriptstyle.006$ $0.011\scriptscriptstyle\pm\scriptstyle.002$ $0.011\scriptscriptstyle\pm\scriptstyle.002$ $0.027\scriptscriptstyle\pm\scriptstyle.004$ $0.034\scriptscriptstyle\pm\scriptstyle.002$ $0.027\scriptscriptstyle\pm\scriptstyle.006$ $0.029\scriptscriptstyle\pm\scriptstyle.003$ $0.014\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.010\scriptscriptstyle\pm\scriptstyle.005}$ $0.035\scriptscriptstyle\pm\scriptstyle.007$ $0.013\scriptscriptstyle\pm\scriptstyle.002$
TVAE $0.025\scriptscriptstyle\pm\scriptstyle.005$ $0.034\scriptscriptstyle\pm\scriptstyle.006$ $0.018\scriptscriptstyle\pm\scriptstyle.000$ $0.014\scriptscriptstyle\pm\scriptstyle.002$ $0.026\scriptscriptstyle\pm\scriptstyle.003$ $0.019\scriptscriptstyle\pm\scriptstyle.001$ $0.027\scriptscriptstyle\pm\scriptstyle.003$ $0.020\scriptscriptstyle\pm\scriptstyle.001$ $0.016\scriptscriptstyle\pm\scriptstyle.002$ $0.030\scriptscriptstyle\pm\scriptstyle.006$ $0.050\scriptscriptstyle\pm\scriptstyle.009$ $0.028\scriptscriptstyle\pm\scriptstyle.004$
CTGAN $0.015\scriptscriptstyle\pm\scriptstyle.001$ $0.017\scriptscriptstyle\pm\scriptstyle.001$ $0.051\scriptscriptstyle\pm\scriptstyle.002$ $0.037\scriptscriptstyle\pm\scriptstyle.002$ $0.047\scriptscriptstyle\pm\scriptstyle.006$ $030$003 $0.037\scriptscriptstyle\pm\scriptstyle.004$ $0.033\scriptscriptstyle\pm\scriptstyle.004$ $0.036\scriptscriptstyle\pm\scriptstyle.005$ $0.018\scriptscriptstyle\pm\scriptstyle.003$ $0.055\scriptscriptstyle\pm\scriptstyle.006$ $0.016\scriptscriptstyle\pm\scriptstyle.003$
TabDDPM $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.008\scriptscriptstyle\pm\scriptstyle.001}$ $0.012\scriptscriptstyle\pm\scriptstyle.001$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.021\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.006\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.017\scriptscriptstyle\pm\scriptstyle.003}$ $\mathbf{0.008\scriptscriptstyle\pm\scriptstyle.001}$ $\mathbf{0.011\scriptscriptstyle\pm\scriptstyle.003}$ $0.017\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.027\scriptscriptstyle\pm\scriptstyle.007}$ $\mathbf{0.010\scriptscriptstyle\pm\scriptstyle.001}$
GReaT $0.014\scriptscriptstyle\pm\scriptstyle.002$ $0.014\scriptscriptstyle\pm\scriptstyle.004$ $0.049\scriptscriptstyle\pm\scriptstyle.002$ $0.029\scriptscriptstyle\pm\scriptstyle.003$ $0.028\scriptscriptstyle\pm\scriptstyle.003$ $0.011\scriptscriptstyle\pm\scriptstyle.001$ $030$003 $0.014\scriptscriptstyle\pm\scriptstyle.001$ $0.019\scriptscriptstyle\pm\scriptstyle.003$ $0.041\scriptscriptstyle\pm\scriptstyle.007$ $0.013\scriptscriptstyle\pm\scriptstyle.001$
PGM ($\epsilon=1$) $0.071\scriptscriptstyle\pm\scriptstyle.014$ $0.052\scriptscriptstyle\pm\scriptstyle.017$ $\mathbf{0.012\scriptscriptstyle\pm\scriptstyle.001}$ $0.036\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.045\scriptscriptstyle\pm\scriptstyle.002}$ $\mathbf{0.037\scriptscriptstyle\pm\scriptstyle.002}$ $0.043\scriptscriptstyle\pm\scriptstyle.005$ $\mathbf{0.050\scriptscriptstyle\pm\scriptstyle.008}$ $\mathbf{0.041\scriptscriptstyle\pm\scriptstyle.003}$ $0.043\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.033\scriptscriptstyle\pm\scriptstyle.004}$ $\mathbf{0.045\scriptscriptstyle\pm\scriptstyle.004}$
PrivSyn ($\epsilon=1$) $\mathbf{0.010\scriptscriptstyle\pm\scriptstyle.001}$ $0.027\scriptscriptstyle\pm\scriptstyle.007$ $0.016\scriptscriptstyle\pm\scriptstyle.002$ $\mathbf{0.025\scriptscriptstyle\pm\scriptstyle.003}$ $0$006 $0.048\scriptscriptstyle\pm\scriptstyle.004$ $0.060\scriptscriptstyle\pm\scriptstyle.004$ $0.095\scriptscriptstyle\pm\scriptstyle.002$ $0.051\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.027\scriptscriptstyle\pm\scriptstyle.005}$ $0.062\scriptscriptstyle\pm\scriptstyle.007$ $0.064\scriptscriptstyle\pm\scriptstyle.008$
PATE-GAN ($\epsilon=1$) $0.028\scriptscriptstyle\pm\scriptstyle.004$ $\mathbf{0.024\scriptscriptstyle\pm\scriptstyle.002}$ $0.117\scriptscriptstyle\pm\scriptstyle.009$ $0.058\scriptscriptstyle\pm\scriptstyle.005$ $0.088\scriptscriptstyle\pm\scriptstyle.009$ $0.191\scriptscriptstyle\pm\scriptstyle.017$ $\mathbf{0.037\scriptscriptstyle\pm\scriptstyle.001}$ $0.150\scriptscriptstyle\pm\scriptstyle.023$ $0.223\scriptscriptstyle\pm\scriptstyle.032$ $0.029\scriptscriptstyle\pm\scriptstyle.003$ $0.138\scriptscriptstyle\pm\scriptstyle.011$ $0.158\scriptscriptstyle\pm\scriptstyle.013$
TableDiffusion ($\epsilon=1$) $0.057\scriptscriptstyle\pm\scriptstyle.006$ $0.054\scriptscriptstyle\pm\scriptstyle.005$ $0.071\scriptscriptstyle\pm\scriptstyle.007$ $0.074\scriptscriptstyle\pm\scriptstyle.011$ $0.119\scriptscriptstyle\pm\scriptstyle.009$ $0.052\scriptscriptstyle\pm\scriptstyle.007$ $0.108\scriptscriptstyle\pm\scriptstyle.010$ $0.071\scriptscriptstyle\pm\scriptstyle.012$ $0.085\scriptscriptstyle\pm\scriptstyle.010$ $0.050\scriptscriptstyle\pm\scriptstyle.003$ $0.195\scriptscriptstyle\pm\scriptstyle.011$ $0.048\scriptscriptstyle\pm\scriptstyle.006$
Appendix E Misleading conclusion
We find that some statements in data synthesis may be misleading or even incorrect due to the lack of principled evaluations. We highlight some of them as below:
  • In [32], authors show that the machine learning performance on TabDDPM is even better than that on real data, which implies that synthetic data can be a perfect (even better) substitute for real data. However, this statement may be incorrect due to inadequate model tuning and improper data shuffling practices. Our evaluations show that even simple models, such as linear regression, can achieve better performance on real data than on high-quality synthetic data (i.e., generated by TabDDPM).
  • Some studies [32, 34, 31, 30] prioritize machine learning efficacy as the primary (if not only) fidelity evaluation metric. This approach is problematic because data synthesis can be biased to label attributes, and a high machine learning efficacy score does not necessarily equate to high fidelity in synthetic data.
  • Extensive studies [71, 34, 31, 32, 68] use Distance to Closest Records (DCR) to evaluate the privacy of synthetic data and assert their models are safe. However, in this paper we show that DCR fails to serve as an adequate measure of privacy. We also provide evidence that many recently introduced HP methods exhibit significant risks of membership disclosure, which are often ignored by HP synthesis algorithms.